### v1

# Structural Alignment Improves Graph Test-Time Adaptation

Abstract
Graph-based learning has achieved remarkable success in domains ranging from recommendation to fraud detection and particle physics by effectively capturing underlying interaction patterns. However, it often struggles to generalize when distribution shifts occur, particularly those involving changes in network connectivity or interaction patterns. Existing approaches designed to mitigate such shifts typically require retraining with full access to source data, rendering them infeasible under strict computational or privacy constraints. To address this limitation, we propose a test-time structural alignment (TSA) algorithm for Graph Test-Time Adaptation (GTTA), a novel method that aligns graph structures during inference without revisiting the source domain. Built upon a theoretically grounded treatment of graph data distribution shifts, TSA integrates three key strategies: an uncertainty-aware neighborhood weighting that accommodates structure shifts, an adaptive balancing of self-node and neighborhood-aggregated representations driven by node representations’ signal-to-noise ratio, and a decision boundary refinement that corrects remaining label and feature shifts. Extensive experiments on synthetic and real-world datasets demonstrate that TSA can consistently outperform both non-graph TTA methods and state-of-the-art GTTA baselines.
Machine Learning, ICML
Figure 1: Example of the distribution shifts of neighborhood information due to different experimental conditions in the LHC experiments. The goal of pileup mitigation is to detect LC neutral particles. The particles are modeled by kNN graphs (dashed circles in the figure) to use nearby particles for drawing inference. The model is trained on ${\mathcal{G}}^{\mathcal{S}}$ but needs to generalize to ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$ over time. The inferred nodes within the circles are the LC neutral particles, but their neighborhood node label ratio changes in ${\mathcal{G}}^{{\mathcal{T}}_{1}}$, ${\mathcal{G}}^{{\mathcal{T}}_{2}}$. In ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ the homophily ratio changes as one of the neighbor node is an LC neutral particle. Both cases represent neighborhood shift which this work aims to address. A more formal definition of these shifts is discussed in Sec. 3.
1 Introduction
Graph-based methods have become indispensable in handling structured data across a wide range of real-world applications (Duvenaud et al., 2015; Bronstein et al., 2017; Zhang et al., 2019a; Stokes et al., 2020), achieving significant success when training and testing data originate from similar distributions. However, these methods struggle to generalize to test-time data in a different domain, where variations in time, location, or experimental conditions result in distinct graph connection patterns. Although some literature on graph domain adaptation (GDA) (Wu et al., 2020; You et al., 2023; Zhu et al., 2021; Liu et al., 2023) seeks to bridge these gaps by aligning labeled source distributions with target distributions, such approaches are often infeasible in practice due to the cost of retraining and the limited availability of source data in the test time.
Figure 2: TSA utilizes neighborhood alignment and SNR-inspired adjustment to mitigate neighborhood information discrepancy. It further adjusts the decision boundary to get refined predictions $\hat{y}$. The refined soft pseudo-labels are used to estimate parameter $\bm{\gamma}$ for neighborhood alignment and to optimize $\alpha$ for combining self representations with neighborhood aggregated representations.
Real-world applications may operate under constraints such as lightweight computation, limited storage, and strict privacy requirements, which make reprocessing of source data generally infeasible (Wang et al., 2021c; Wu et al., 2021). For example, in particle tracking based on graph learning at the Large Hadron Collider (LHC) (Shlomi et al., 2020; Highfield, 2008; Miao et al., 2024), the connections between particles differ as experimental conditions change over time. The tracker is expected to be capable of adapting on-the-fly (Li et al., 2023; Komiske et al., 2017). Similarly, in fraud detection for financial networks  (Clements et al., 2020; Wang et al., 2021b), privacy-sensitive training data may not be accessed during testing, while transaction patterns may change between different regions  (Wang et al., 2019; Dou et al., 2020). These scenarios indicate the need for graph test-time domain adaptation (GTTA), which enables models to adjust to new test domains without re-accessing the source domain or incurring the overhead of full retraining.
Accurate node classification in graph-structured data relies heavily on effectively leveraging neighborhood information. Graph neural networks (GNNs) (Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017; Hamilton, 2020) have achieved significant success across a wide range of applications by utilizing this neighborhood context, yet they remain sensitive to distribution shifts in node neighborhoods (Gui et al., 2022; Ji et al., 2023). As shown in Fig. 1, shifts in LHC data can lead to distinct neighborhood connections among center particles in the same class (neighborhood shift). Existing methods for test-time adaptation (TTA) often focus on adjusting only the classifier’s final layer (Iwasawa & Matsuo, 2021), modifying its outputs (Boudiaf et al., 2022), or adapting normalization statistics (Wang et al., 2021a; Niu et al., 2023), primarily targeting image-based tasks with independently labeled data points. Consequently, they may not adequately address neighborhood shifts for node classification in graphs. While recent works on GTTA attempt to address the structural changes using graph augmentation (Jin et al., 2023), self-supervised learning (Mao et al., 2024). These approaches mainly rely on heuristics and homophily consistency, failing to fully address changes in neighboring node connections in principle.
In this work, we propose a model-agnostic framework with test-time structural alignment, named TSA, as illustrated in Fig. 2, and supported by theoretical analysis. We begin by examining the generalization gap in GTTA for a pretrained source model. Our theoretical findings, further validated by empirical studies on synthetic datasets, indicate the importance of aligning the label distributions of neighboring nodes among the same class of nodes across the source and target domains. In the context of training-time GDA, PairAlign (Liu et al., 2024) employs a neighborhood weighting strategy to recalibrate the influence of neighboring nodes during message aggregation for a similar alignment. However, extending this approach to GTTA necessitates our more nuanced investigation, leading to the following contributions:
Firstly, the weight assignment relies on knowledge of node labels, which are unavailable in the target graph under GTTA. Assigning weights based on pseudo labels may, in fact, degrade performance. To mitigate this, TSA introduces an uncertainty-aware assignment strategy that aligns only node pairs with more reliable test-time pseudo labels.
Additionally, TSA optimizes the test-time combination of self-node representations and neighborhood-aggregated representations based on their signal-to-noise ratio (SNR). For instance, when the target graph is sparse, TSA reduces reliance on neighborhood messages due to their high variance and low SNR. Conversely, as layer depth increases, it assigns greater weight to self-node representations, which become progressively denoised.
Lastly, TSA integrates non-graph TTA methods to refine the decision boundary, mitigating mismatches caused by additional label and feature shifts once neighborhood shift has been addressed.
We conduct extensive experiments on synthetic and four real-world datasets, including those from high-energy physics and citation networks, with multiple GNN backbones, to demonstrate the effectiveness of TSA. TSA outperforms non-graph TTA baselines by up to 12% on synthetic datasets and surpasses all non-graph baselines. Compared to existing GTTA baselines, TSA achieves an average of 10% improvement on all real-world datasets.
2 Preliminaries and Related Works
2.1 Notations and Problem Setup
We use upper-case letters, such as $Y$ to represent random variables, lower-case letters, such as $y$ to represent their realization. The calligraphic letters, such as ${\mathcal{Y}}$ denote the domain of random variables. We use bold capital such as ${\mathbf{Y}}$ to represent the vectorized corresponds, i.e., a collections of random variables. The probability distribution of a random variable $Y$ for a realization is expressed as $\mathbb{P}(Y=y)$.
Graph Neural Networks (GNNs).
We let ${\mathcal{G}}=({\mathcal{V}},{\mathcal{E}})$ denote an undirected and unweighted graph with the symmetric adjacency matrix ${\bm{A}}\in\mathbb{R}^{N\times N}$ and the node feature matrix ${\bm{X}}=[x_{1},\dots,x_{N}]^{T}$. GNNs utilize neighborhood information by encoding ${\bm{A}}$ and ${\bm{X}}$ into node representations $\{h_{v}^{(k)},v\in\mathcal{N}_{u}\}$. With $h_{u}^{(1)}=x_{u}$, the message passing in standard GNNs for node $v$ and each layer $k\in[L]\coloneqq\{1,\dots,L\}$ can be written as
$h_{u}^{(k+1)}=\text{UPT}(h_{u}^{(k)},\text{AGG}(\{h_{v}^{(k)},v\in\mathcal{N}_{u}\}))$ (1)
where $\mathcal{N}_{u}$ denotes the set of neighbors of node $u$, which $|\mathcal{N}_{u}|$ represent the node degree $d_{u}$. The AGG function aggregates messages from the neighbors, and the UPT function updates the node representations.
Graph Test-Time Adaptation (GTTA).
Assume the model consists of a GNN feature encoder $\phi:{\mathcal{X}}\rightarrow{\mathcal{H}}$ and a classifier $g:{\mathcal{H}}\rightarrow{\mathcal{Y}}$. The model is trained on source graph ${\mathcal{G}}^{{\mathcal{S}}}=({\mathcal{V}}^{{\mathcal{S}}},{\mathcal{E}}^{{\mathcal{S}}})$ with node labels $\mathbf{y}^{\mathcal{S}}$ and the goal is to enhance model performance on test graph ${\mathcal{G}}^{{\mathcal{T}}}=({\mathcal{V}}^{{\mathcal{T}}},{\mathcal{E}}^{{\mathcal{T}}})$ with distribution shifts that will be defined in Sec. 3.1. For node classification tasks, we aim to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})$ without accessing ${\mathcal{G}}^{{\mathcal{S}}}$.
2.2 Related Works
Test-time Adaptation Test-time training (Sun et al., 2020; Liu et al., 2021; Bartler et al., 2022) adapts the source model to the target domain but requires to first add a customized self-supervised losses in model pretraining. In contrast, our setup falls into the category of fully test-time adaptation (Wang et al., 2021a; Liang et al., 2024), where we do not alter the model training pipeline. Tent (Wang et al., 2021a) adapts the batch normalization parameters by minimizing the entropy, which motivated following-up studies on normalization layer adaptation  (Gong et al., 2022; Zhao et al., 2023; Lim et al., 2023; Niu et al., 2023). Some TTA works directly modify the classifier’s prediction, such as LAME (Boudiaf et al., 2022) that directly refines the model’s soft prediction by applying a regularization in the feature space, T3A (Iwasawa & Matsuo, 2021) that classifies test data based on the distance to the pseudo-prototypes derived from pseudo-labels, TAST (Jang et al., 2022) and PROGRAM (Sun et al., 2024) that extend T3A through constructing more reliable prototype graphs. However, the above methods are designed for image-based applications and cannot handle the shifts in neighborhood information of graph data.
Graph Test-time Adaptation Studies on GTTA are in two categories - node and graph classification. Graph classification problems can treat each graph as an i.i.d. input, allowing more direct extension of image-based TTA techniques to graphs (Chen et al., 2022; Wang et al., 2022). Our work focuses on node classification. GTrans (Jin et al., 2023) proposes to augment the target graph at the test time by optimizing a contrastive loss by generating positive views from DropEdge (Rong et al., 2020) and negative samples from the features of the shuffling nodes (Veličković et al., 2019). GraphPatcher (Ju et al., 2024) learns to generate virtual neighbors to improve low-degree nodes classification. SOGA (Mao et al., 2024) designs a self-supervised loss that relies on mutual information maximization and homophily assumption. These works are mostly built upon heuristics and may not address structure shifts in principle.
3 Test Error Analysis
In this section, we characterize the generalization error between source and target graphs and explicitly attribute it to three different kinds of shifts: label shift, feature shift, as well as neighborhood shift. Motivated by our theoretical analysis on the generalization error, we then propose TSA algorithm to minimize the across-graph generalization error.
(a) Reliability of $\gamma$ (b) Source (c) Nbr. Align (d) SNR $\alpha$
Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 3: (a) Comparison of neighborhood alignment with $\bm{\gamma}$ from model prediction and Oracle on the CSBM graphs (Deshpande et al., 2018). The top (or bottom) subfigures represents the assignment under neighbor shift (or neighbor shift plus label shift, respectively). Nodes are grouped by the entropy of their soft pseudo labels and the y axis shows the accuracy after assigning $\bm{\gamma}$. Ideally, a correct assignment (red) would lead to near-perfect accuracy. However, the assignment based on pseudo labels is far from optimal (blue). From (b) to (c), the figure with the t-SNE visualization of node representations indicates a model trained from the source domain (CSBM with a label distribution $[0.1,0.3,0.6]$) to the target domain (CSBM with a label distribution $[0.3,0.3,0.3]$). The color of the nodes represents the ground-truth labels. The top subfigures of (b) and (c) show the output given by the GNN encoder while the bottom subfigures show the classifier decision boundaries. (d) Analysis of SNR adjustment $\alpha$ with respect to different layers and node degrees on MAG. Detail discussed in Sec. 5.3.
3.1 Distribution Shifts on Graphs
Distribution shifts on graphs were formally studied in previous GDA works (Wu et al., 2020; Liao et al., 2021; Zhu et al., 2021; Wu et al., 2022; You et al., 2023; Zhu et al., 2023). Following their definition, we categorize shifts in graphs into two types: feature shift and structure shift. For simplicity, our analysis is based on a data generation process: ${\mathbf{X}}\leftarrow{\mathbf{Y}}\rightarrow{\mathbf{A}}$, where graph structure and node features are both conditioned on node labels.
Definition 3.1.
(Feature Shift). Assume node features $x_{u},u\in{\mathcal{V}}$ are i.i.d sampled given labels $y_{u}$, then we have $\mathbb{P}({\mathbf{X}}|{\mathbf{Y}})=\prod_{u\in{\mathcal{V}}}\mathbb{P}(X_{u}|Y_{u})$. We then define the feature shift as $\mathbb{P}^{\mathcal{S}}(X_{u}|Y_{u})\neq\mathbb{P}^{\mathcal{T}}(X_{u}|Y_{u})$.
Definition 3.2.
(Structure Shift). As graph structure involves the connecting pattern between labels, we consider the joint distribution of the adjacency matrix and labels $\mathbb{P}({\mathbf{A}},{\mathbf{Y}})$, where Structure shift, denoted by $\mathbb{P}^{\mathcal{S}}({\mathbf{A}},{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}},{\mathbf{Y}})$, can be decomposed into as conditional structure shift (CSS) $\mathbb{P}^{\mathcal{S}}({\mathbf{A}}|{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}}|{\mathbf{Y}})$ and label shift (LS) $\mathbb{P}^{\mathcal{S}}({\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{Y}})$.
3.2 Theoretical Analysis
Let $\varepsilon^{\mathcal{S}}(g\circ\phi)$ denote the error of a pretrained GNN on the source domain and $\varepsilon^{\mathcal{T}}(g\circ\phi)$ the error of the model when applied to a test graph in the target domain. Inspired by Tachet des Combes et al. (2020), we provide an upper bound on the error gap between source and target domains, showing how a pretrained GNN (e.g., ERM) can be influenced. We denote the balanced error rate of a pretrained node predictor $\hat{Y}_{u}$ on the source domain as $\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\coloneq\max_{j\in\mathcal{Y}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}\neq Y_{u}|Y_{u}=j).$
Theorem 3.3 (Error Decomposition Theorem).
Suppose ${\mathcal{G}}^{{\mathcal{S}}}$ and ${\mathcal{G}}^{{\mathcal{T}}}$, we can decouple both graphs into independent ego-networks (center nodes and 1-hop neighbors). For any classifier $g$ with a mean pooling GNN encoder $\phi$ in node classification tasks, we have the following upper bound for on the error gap between source and target under feature shift and structure shift:
$\lvert\varepsilon^{\mathcal{S}}(g\circ\phi)-\varepsilon^{\mathcal{T}}(g\circ\phi)\rvert$
$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2\text{TV}(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}$
$+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Neighborhood shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature shift}}$
where $TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))$ is the total variation distance between the source and target label distributions. and $\Delta_{CE}$ is the error gap that exists if and only if feature shift exists.
Our bound aligns with the findings of Liu et al. (2024), which highlight the impact of neighborhood shift, label shift, and feature shift on generalization to the target graph. We extend this understanding by deriving an explicit error bound. Notably, neighborhood shift is reduced from conditional structure shift given the assumptions in Thm.3.3.
4 Test-Time Structural Alignment (TSA)
Motivated by our theoretical analysis, we propose TSA to address GTTA in a principled way. To address neighborhood shift, TSA first conducts neighborhood alignment via weighted aggregation and then properly balances self and neighboring node information combination based on the SNRs of these representations. Lastly, as a generic approach for neighborhood shift, TSA can be combined with non-graph TTA methods to refine the decision boundary to address the remaining feature and label shift.
4.1 Neighborhood Alignment
Neighborhood shift alters the node label ratios in the aggregated neighborhood information, causing the shift in the GNN-encoded representations. In training-time GDA, PairAlign (Liu et al., 2024) leverages such a technique by assigning edge weights to align the distribution of the source neighborhood with the target domain.
Inspired by this idea, for GTTA, we aim to achieve a similar goal but in a different direction. Based on Theorem 3.3, we align the target neighborhood distribution with the source domain, leveraging the fact that the pre-trained model is optimized for the source distribution. Specifically, the neighborhood distribution determines the ratio of messages passed from a neighboring class $j$ to center class $i$. To adjust for distributional differences, this ratio can be rescaled by assigning weights to edges from class $j$ to class $i$, effectively acting as an up/down-sampling mechanism during message aggregation. To ensure that message aggregation in the target domain aligns with the expected behavior in the source domain, TSA incorporates the following adjustment:
Definition 4.1.
Let $\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})>0,\forall i,j\in{\mathcal{Y}}$, we have $\bm{\gamma}\in\mathbb{R}^{|{\mathcal{Y}}|\times|{\mathcal{Y}}|}$ as:
$[\bm{\gamma}]_{i,j}=\frac{\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})},\>\forall i,j\in{\mathcal{Y}}$ (2)
To estimate $\bm{\gamma}$, we assume that the source summary statistics $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ are recorded and available at test time; otherwise, alignment would not be feasible. Storing $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ incurs minimal cost, as it is merely an $|{\mathcal{Y}}|\times|{\mathcal{Y}}|$ matrix. Beyond this, no additional information from the source domain is required. For $\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$, we estimate it based on target pseudo labels. Note that PairAlign (Liu et al., 2024) enhances estimation accuracy that relies pseudo labels (Liu et al., 2023) by leveraging a least-squares constrained optimization. However, in GTTA, the absence of source graphs and the potential need for real-time adaptation render this approach impractical.
While it seems direct to apply a similar alignment strategy by using the inverse of the ratio employed in PairAlign, assigning $\gamma$ requires knowledge of node labels, which is missing in the targe graph, making it challenging in GTTA.
Reliable assignment of $\bm{\gamma}$.
The ratio $\gamma$ should be assigned to edge weights based on the label pairs of the central and neighboring nodes. In training-time GDA, this assignment is straightforward as it relies on the ground-truth labels of the source graph. However, in GTTA, this information is unavailable. A naive approach is to use target pseudo labels, but this often results in significant mismatches. TSA addresses this by quantifying the uncertainty of target pseudo labels (Zhang et al., 2019b; Stadler et al., 2021; Hsu et al., 2022). In particular, TSA assigns $[\bm{\gamma}]_{i,j}$ only to node pairs $v\rightarrow u$ where both of their soft target pseudo labels $\hat{y}$ have low entropy $H(\hat{y})=-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$. Here, $\rho_{1}$ is a hyperparameter and $\ln(|{\mathcal{Y}}|)$ is the maximum entropy in $|{\mathcal{Y}}|$ class prediction. In Fig.3 (a), nodes with low-entropy soft predictions are more reliable, resulting in higher accuracy after the assignment of $\bm{\gamma}$.
Algorithm 1 Test-Time Structural Alignment (TSA)
1:  Input: A GNN $\phi$ and a classifier $g$ pretrained on source graph ${\mathcal{G}}^{{\mathcal{S}}}$; Test-time target graph ${\mathcal{G}}^{{\mathcal{T}}}$; Source statistics $\mathbb{P}^{\mathcal{S}}(Y_{v}|Y_{u},v\in\mathcal{N}_{u})\in\mathbb{R}_{+}^{|{\mathcal{Y}}|\times|{\mathcal{Y}}|}$
2:  Initialize $b^{(k)}\leftarrow 1$ and $\text{MLP}^{(k)}\text{ parameters}\leftarrow 0$ for the $k$-th layer.
3:  Perform boundary refinement based on embeddings from $g\circ\phi({\mathcal{G}}^{{\mathcal{T}}})$ and get soft pseudo-labels $\hat{y}$
4:  Get ${\mathcal{G}}^{{\mathcal{T}}}_{new}$ by assigning edge weights:
5:     Compute $\gamma$ using $\hat{y}$ via Eq. 2
6:     Assign $\gamma$ only if node pairs $H(\hat{y})\leq\rho_{1}\cdot\ln(|\mathcal{Y}|)$
7:     Assign the parameterized $\alpha$ via Eq. 3
8:  Update $\alpha$’s parameters $b^{(k)}$ and $\text{MLP}^{(k)}$ via Eq. 4
9:  return $\hat{y}_{\mathrm{final}}$ after another boundary refinement
4.2 SNR-Inspired Adjustment
Building on neighborhood alignment, we further optimize the signal-to-noise ratio (SNR) of node representations to enhance performance. Specifically, SNR is characterized by the ratio of inter-class representation distances to intra-class representation distances. A higher SNR indicates more informative and well-separated representations, which benefit classification.
Optimizing SNR complements the neighborhood alignment approach. Even if neighborhood label distributions are perfectly aligned, variations in neighbor set sizes between the source and target graphs can impact the SNR of aggregated neighboring node representations. Consequently, the combination of these aggregated representations with self-node representations in a typical GNN pipeline (Eq. 1) should be adjusted accordingly across source and target domains, particularly when the source and target graphs exhibit very different degree distributions. Furthermore, the SNR of node self-representations may vary across GNN layers, as deeper layers generally reduce variance. As a result, node self-representations in deeper layers tend to have higher SNR and should be assigned greater emphasis.
To implement the SNR-inspired adjustment, we introduce a parameter to perform the weighted combination of self-node representations and neighborhood-aggregated representations at each layer, adapting to node degrees as follows:
Definition 4.2.
Let $\tilde{d_{u}}=\left(\frac{\ln(d_{u}+1)}{\ln(\max_{v\in\mathcal{V}}d_{v}+1)}\right)$ denote log-normalized degree of node $u$ and let $\text{MLP}^{(k)}$ and $b^{(k)}$ to be learnable parameters for adjusting $k$-th layer combination, define the weights for combination $\alpha\in\mathbb{R}^{L}$ as:
$[\alpha]_{k}=\sigma(\text{MLP}^{(k)}(\tilde{d_{u}}))-0.5+b^{(k)},\>\forall k\in[L]$ (3)
where $\sigma$ is the sigmoid function. During initialization $[\alpha]_{k}$ is set to $1$. Degree values are taken in the logarithmic domain to handle their often long-tailed distribution.
Combined with $\bm{\gamma}$, $[\alpha]_{k}\cdot[\bm{\gamma}]_{i,j}$ is used to reweight the GNN message for non-self-loop node pairs, adjusting the influence from the node with a highly certain pseudo label $j$ to the node with a highly certain pseudo label $i$.
Remark. Neighborhood alignment alone does not address potential shifts in SNR. This is because the alignment approach, inspired by Thm. 3.3, focuses solely on expectation-based (first-order statistical) performance, whereas SNR also incorporates variance (second-order statistics). Thus, these two aspects complement each other.
4.3 Decision Boundary Refinement
In GTTA, label shift can result in a mismatch of the decision boundary, even after addressing neighborhood shift and obtaining high-SNR node representations. This is illustrated in Fig. 3 (b) and (c).
A straightforward approach to refining the decision boundary at test time is to adjust the classifier’s batch normalization parameters (TENT (Wang et al., 2021a)) or directly modify its output (T3A (Iwasawa & Matsuo, 2021) and LAME  (Boudiaf et al., 2022)). We integrate these techniques into our framework for two folds: (1) their refined pseudo-labels provide a more reliable assignment of $\bm{\gamma}$ and can supervise the update of SNR adjustment. (2) Reciprocally, better alignment of neighborhood information can further refine the decision boundary.
4.4 TSA Overview.
Note that $\bm{\gamma}$ is obtained from the initial pseudo labels. Only the parameters in $\text{MLP}^{(k)}$ and $b^{(k)}$ for $\alpha$ estimation in Eq. 3 need to be optimized in the test time according to Eq. 4.
${\mathcal{L}}_{CE}=\frac{1}{|{\mathcal{V}}^{{\mathcal{T}}}|}\sum_{u\in{\mathcal{V}}^{{\mathcal{T}}}}\text{cross-entropy}(y_{u}^{\prime},\hat{y}_{u})$ (4)
where $y_{u}^{\prime}$ is the hard pseudo label refined by procedure in Sec. 4.3 and $\hat{y}_{u}$ is the soft prediction from the original model. After updating $\alpha$, TSA makes predictions on the newly weighted graphs and then further adapts the boundaries as described in Sec. 4.3. Our proposed algorithm is summarized in Alg. 1.
Comparison to AdaRC (Bao et al., 2024) AdaRC is a recent work on GTTA that also integrates non-graph TTA methods with an approach for addressing structure shifts. However, it considers only degree shift and homophily shift, where homophily shift is merely a special case of neighborhood shift in our context. Consequently, AdaRC does not introduce the parameter $\bm{\gamma}$ and, therefore, lacks the ability to properly align neighborhood aggregated representations when the neighboring label distribution shifts.
5 Experiments
Table 1: MAG results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
[…]
Table 2: Synthetic CSBM results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model. First six: imbalanced source training. Last two: balanced source training.
Nbr. Shift Nbr.+ SNR Shift Struct. Shift (Imbal.$\rightarrow$ Bal.) Struct. Shift (Bal.$\rightarrow$ Imbal.)
ERM 82.70$\pm$4.45 61.11$\pm$10.81 77.03$\pm$3.99 61.93$\pm$6.44 50.41$\pm$4.88 39.12$\pm$4.71 68.27$\pm$5.00 61.39$\pm$1.89
GTrans 86.67$\pm$3.59 72.37$\pm$4.06 79.55$\pm$1.23 68.69$\pm$3.27 59.24$\pm$2.12 47.42$\pm$3.73 79.29$\pm$2.71 66.77$\pm$2.52
SOGA 86.09$\pm$3.89 70.39$\pm$7.96 79.75$\pm$3.20 69.00$\pm$5.28 56.60$\pm$3.88 44.09$\pm$5.32 73.52$\pm$5.30 63.76$\pm$3.45
TENT 87.48$\pm$2.86 77.14$\pm$4.64 81.04$\pm$2.72 72.51$\pm$3.39 76.48$\pm$6.21 59.12$\pm$5.06 77.21$\pm$5.53 62.36$\pm$6.83
LAME 83.96$\pm$5.35 61.44$\pm$11.33 77.56$\pm$4.92 62.58$\pm$7.24 50.33$\pm$4.78 39.05$\pm$4.67 68.33$\pm$5.02 61.26$\pm$1.97
T3A 77.05$\pm$7.10 59.83$\pm$10.50 71.44$\pm$6.11 56.56$\pm$7.50 48.13$\pm$5.64 38.19$\pm$3.72 68.50$\pm$4.81 61.63$\pm$1.81
TSA-TENT 88.78$\pm$1.37 80.51$\pm$2.39 83.19$\pm$1.46 76.41$\pm$1.25 88.68$\pm$4.99 66.25$\pm$7.75 81.20$\pm$8.18 70.15$\pm$2.30
TSA-LAME 88.96$\pm$1.66 80.02$\pm$5.44 83.51$\pm$0.55 79.56$\pm$1.82 65.09$\pm$2.34 52.90$\pm$6.11 82.20$\pm$5.17 63.15$\pm$2.58
TSA-T3A 89.96$\pm$1.33 81.08$\pm$2.73 84.23$\pm$1.24 76.89$\pm$2.02 65.59$\pm$2.57 52.34$\pm$7.19 82.55$\pm$5.06 64.88$\pm$3.30
Table 3: Pileup results (f1-scores). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
[…]
ERM 57.98$\pm$0.66 65.40$\pm$2.17 47.66$\pm$1.47 67.81$\pm$1.70 19.42$\pm$2.59 57.49$\pm$3.02 69.35$\pm$0.81 67.90$\pm$0.46
GTrans 57.37$\pm$1.49 63.66$\pm$2.43 48.13$\pm$2.11 65.74$\pm$1.95 28.41$\pm$4.01 57.65$\pm$1.97 69.17$\pm$0.82 67.37$\pm$0.56
SOGA 60.13$\pm$0.75 69.71$\pm$0.92 51.68$\pm$0.83 67.47$\pm$1.42 37.16$\pm$1.15 56.59$\pm$3.86 70.83$\pm$0.66 68.84$\pm$0.97
TENT 58.25$\pm$1.85 58.73$\pm$4.80 48.27$\pm$2.58 58.41$\pm$7.89 30.59$\pm$3.07 0.04$\pm$0.08 68.72$\pm$0.48 68.00$\pm$0.60
LAME 57.77$\pm$0.66 66.29$\pm$2.42 47.31$\pm$1.58 68.24$\pm$1.61 11.96$\pm$3.08 58.81$\pm$2.00 69.25$\pm$0.52 67.94$\pm$0.66
T3A 58.74$\pm$1.09 70.02$\pm$2.33 49.61$\pm$1.03 70.85$\pm$0.64 30.33$\pm$3.66 54.79$\pm$1.52 69.45$\pm$0.97 68.45$\pm$0.61
TSA-TENT 58.27$\pm$1.84 60.11$\pm$3.67 48.44$\pm$2.67 58.79$\pm$7.01 36.22$\pm$2.04 6.13$\pm$5.93 69.26$\pm$0.54 68.80$\pm$0.78
TSA-LAME 60.76$\pm$0.71 66.94$\pm$1.62 50.72$\pm$1.49 68.53$\pm$1.55 36.43$\pm$2.57 59.36$\pm$2.53 69.85$\pm$0.66 68.31$\pm$0.54
TSA-T3A 61.03$\pm$1.19 70.42$\pm$1.66 52.39$\pm$0.86 70.92$\pm$0.64 37.33$\pm$2.59 57.73$\pm$1.98 69.73$\pm$1.09 68.75$\pm$0.66
Table 4: Arxiv and DBLP/ACM no balanced training (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
1950-2007 1950-2009 1950-2011 DBLP & ACM
[…]
ERM 41.04$\pm$0.50 40.48$\pm$1.45 44.80$\pm$1.96 42.38$\pm$3.64 53.40$\pm$1.14 51.68$\pm$1.76 28.95$\pm$4.50 52.45$\pm$6.81
GTrans 40.92$\pm$0.32 40.25$\pm$1.69 45.31$\pm$1.99 43.83$\pm$3.15 53.68$\pm$0.95 52.57$\pm$1.23 34.47$\pm$2.99 49.68$\pm$5.44
SOGA 34.11$\pm$2.91 28.94$\pm$4.68 41.59$\pm$2.03 39.61$\pm$3.22 50.12$\pm$1.38 43.03$\pm$3.70 35.74$\pm$5.20 58.59$\pm$8.35
TENT 40.77$\pm$0.32 39.58$\pm$1.30 45.74$\pm$1.26 43.98$\pm$1.90 54.45$\pm$1.03 53.19$\pm$1.31 36.59$\pm$4.26 53.38$\pm$6.60
LAME 40.91$\pm$0.88 41.02$\pm$1.81 45.13$\pm$2.49 43.23$\pm$4.42 53.63$\pm$1.21 52.07$\pm$1.62 28.02$\pm$5.08 52.74$\pm$7.08
T3A 39.57$\pm$1.06 38.98$\pm$1.78 43.34$\pm$1.51 41.17$\pm$2.88 50.10$\pm$1.70 48.00$\pm$2.82 35.29$\pm$5.58 52.50$\pm$6.83
TSA-TENT 40.92$\pm$0.34 39.78$\pm$1.21 46.68$\pm$1.24 45.08$\pm$1.71 54.78$\pm$0.80 53.61$\pm$1.24 37.06$\pm$3.93 54.06$\pm$6.63
TSA-LAME 41.34$\pm$0.83 41.23$\pm$1.70 45.42$\pm$2.45 43.71$\pm$4.99 54.05$\pm$1.05 52.76$\pm$1.47 28.66$\pm$5.27 52.90$\pm$7.30
TSA-T3A 40.03$\pm$1.03 39.70$\pm$1.49 44.09$\pm$1.53 42.17$\pm$2.91 50.96$\pm$1.68 49.21$\pm$2.67 35.30$\pm$5.57 52.55$\pm$6.92
We evaluate TSA on synthetic datasets and 5 real-world datasets. More discussions and results in experiments can be found in Appendix D.
5.1 Datasets and Baselines
Synthetic Data. We use the CSBM model (Deshpande et al., 2018) to generate three-class datasets, focusing on structure shifts while keeping feature distributions unchanged. Specifically, we evaluate performance under three conditions: (1) neighborhood shift, (2) neighborhood shift plus SNR (induced by degree changes) shift, and (3) neighborhood shift combined with both SNR shift and label shift. This experimental setup is motivated by Thm. 3.3 and the observations in Sec.4.2. For each shift scenario, we examine two levels of severity, with the left column in later Table 2 corresponding to the smaller shift.
MAG (Liu et al., 2024) is a citation network extracted by OGB-MAG (Hu et al., 2020). Distribution shifts arise from partitioning papers into distinct graphs based on their countries of publication. The task is to classify the publication venue of each paper. Our model is pretrained on graphs from the US and China and subsequently adapted to graphs from other countries.
[…]
Arxiv (Hu et al., 2020) is a citation network between all Computer Science (CS) arXiv papers. Distribution shifts originate from different time. Our model is pretrained on the earlier time span 1950 to 2007/ 2009/ 2011 and test on later 2014 to 2016 and 2016 to 2018.
DBLP and ACM (Tang et al., 2008; Wu et al., 2020) are two citation networks. The model is trained on one network and adapted to the other to predict the research topic of a paper (node).
Baselines We compare TSA with six baselines. For non-graph TTA methods, we include TENT (Wang et al., 2021a), LAME (Boudiaf et al., 2022), and T3A (Iwasawa & Matsuo, 2021). Note that TENT is only applied to the classifier, as GNNs typically do not include batch normalization layers. GTrans (Jin et al., 2023), SOGA (Mao et al., 2024), and AdaRC (Bao et al., 2024) are direct comparisons in GTTA. AdaRC is limited to GPRGNN due to its design. We present the results for GraphSAGE (Hamilton et al., 2017) in the main paper, while the results for GPRGNN (Chien et al., 2020) (including AdaRC) are provided in the Appendix.
5.2 Result Analysis
From the MAG dataset results in Table 1, TSA combined with boundary refinement techniques consistently achieves top performance among all baselines. Compared to non-graph TTA baselines, such as TENT vs. TSA-TENT, TSA also provides considerable improvements over the original adaptation results. Notably, the benefits of TSA after neighborhood alignment are more pronounced in scenarios with larger neighborhood shifts, such as adaptation from the US to FR, as reflected in the dataset shift metrics (Table LABEL:table:magstats).
Among all baselines, T3A consistently outperforms as a representative non-gradient-based method that refines the decision boundary, while TENT struggles in most scenarios. This is primarily due to the relatively imbalanced class distribution in the MAG datasets, where evaluation is conducted on the top 19 classes. Since TENT learns feature-wise transformations through entropy optimization on the target graph, it tends to be biased toward the dominant class. Other GTTA methods, such as GTrans, achieve performance comparable to ERM, whereas SOGA sometimes underperforms due to its strong reliance on the graph homophily assumption, which is often weak in sparse graphs with many classes, such as MAG and Arxiv.
The results for the synthetic datasets are presented in Table 2. The first six columns correspond to experiments conducted under imbalanced source training. Compared to real datasets, we observe more significant advantages from TSA. TENT outperforms LAME and T3A, as it directly maximizes the most probable class through entropy minimization. In contrast, LAME incorporates regularization, while T3A relies on distance to prototypes, making it less sensitive to majority-class dominance (Zhang et al., 2022). When combined with TSA, we observe similar results across all three variants, indicating that the primary adaptation ability stems from TSA rather than non-graph TTA methods. While GTTA baselines perform better than ERM, they still fall short of our TSA-based methods. TSA also demonstrates strong performance under additional shifts in SNR caused by degree shifts. Additionally, we consider training a pretrained model on a balanced source domain under the same edge connection probability as the previous two columns. The last two columns demonstrate better performance than the previous two under balanced training, even though they encounter the same shift in label connection. This showcases that dataset imbalance is a severe issue in source training, which aligns with the worst-case error described in Theorem 3.3.
The pileup results in Table 3 demonstrate the effectiveness of TSA-based methods under both neighborhood shift and label shift, with each module contributing in different ways. The neighborhood alignment module provides significant improvements when generalizing from a low PU level to a high PU level, as this scenario is primarily dominated by neighborhood shift. Conversely, when adapting from a high PU level to a low PU level, label shift becomes the dominant factor. In this case, incorporating non-graph TTA techniques is crucial, while neighborhood alignment offers an additional but smaller benefit. Moreover, TENT can be vulnerable when generalizing from an imbalanced source graph to a balanced target graph. This is because source training on high PU graphs suffers from class imbalance, making entropy minimization non-robust due to a mismatched decision boundary when adapting to low PU graphs. The last two columns, which represent the same PU level but different physical signals, show minimal benefits from TSA. This is expected, as these scenarios primarily exhibit feature shifts rather than structure or label shifts.
The results from Arxiv and DBLP/ACM in Table 4 confirm that TSA can be effectively integrated with non-graph TTA methods to mitigate feature shift while also facilitating reasonable adjustments for structure shifts. These citation networks exhibit relatively mild structure shifts, leading to closely comparable performance across all baselines. TENT and TSA-TENT stand out by maximizing the majority classes through entropy minimization. Nonetheless, we still observe meaningful improvements with the addition of TSA when compared to the corresponding baselines, demonstrating its effectiveness in enhancing adaptation.
5.3 Analysis of $\alpha$ for SNR-induced refinement
As defined in Eq.3, $\alpha$ is learned, depending on node degrees and GNN layer depth. The trend of $\alpha$ in Fig. 3 (d) aligns with our expectations: earlier GNN layers require less attention to self-node representations (small $\alpha$), while deeper layers increasingly rely on them. Additionally, the plot reveals a correlation with node degrees, where nodes with higher degrees (and thus higher SNR) place greater weight on neighborhood-aggregated messages. However, the overall effect of different GNN layers appears to be greater than that of varying node degrees.
5.4 Ablation Studies
In Table 6 (Appendix), we examine the effects of neighborhood alignment and SNR-based representation adjustments. The results consistently show that incorporating both modules yields the best performance. In some cases, removing SNR adjustments leads to a greater performance drop compared to removing neighborhood alignment. However, this does not necessarily indicate that neighborhood alignment is less important. Instead, we think this effect arises because the additional backpropagation training for SNR adjustments amplifies their contributions.
6 Discussion and Limitations
In this work, we theoretically analyze the impact of various graph shifts on the generalization gap in GTTA and highlight their empirical effects on test-time performance degradation. Based on these insights, we propose the TSA algorithm, which addresses neighborhood shift through extended and improved neighborhood alignment strategies tailored to the unique challenges of GTTA, and handles label shift via classifier adjustments.
Despite its effectiveness, we observe that TSA ’s benefits can be marginal when feature shift dominates, as evidenced in the last two columns of Pileup, Arxiv, and DBLP/ACM. This is primarily because the neighborhood alignment strategy is not robust in such scenarios. Moreover, our current estimation of $\bm{\gamma}$ heavily depends on the accuracy of refined soft pseudo-labels. While confusion-matrix-based methods (Lipton et al., 2018; Azizzadenesheli et al., 2019; Alexandari et al., 2020) can provide more robust estimations, they often require additional adversarial training, which is not compatible with GTTA constraints. We encourage future work to explore more robust approaches for estimating $\bm{\gamma}$ to further enhance adaptation under severe feature shifts.
7 Broader Impact
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
8 Acknowledgment
H. Hsu, S. Liu, and P. Li are partially supported by NSF awards PHY-2117997, IIS-2239565, IIS-2428777, and CCF-2402816; DOE award DE-FOA-0002785; JPMC faculty awards; and Microsoft Azure Research Credits for Generative AI.
Appendix A Omitted Proofs
A.1 Proof of Theorem 3.3
Lemma A.1.
Assume :-$\mathbb{P}^{\mathcal{M}}\coloneq\alpha_{j}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)+\beta_{j}\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)$ to be a mixture conditional probability of source and target domain. The mixture weight is given by $\alpha_{j}$ and $\beta_{j}$ where $\forall\alpha_{j},\beta_{j}\geq 0$ and $\alpha_{j}+\beta+j=1$, the following upper bound holds:
$\lvert\gamma^{\mathcal{S}}_{u_{j}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\gamma^{\mathcal{T}}_{u_{j}}\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j\rvert$
$\leq\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\cdot\mathbb{P}^{\mathcal{M}}+(\gamma^{\mathcal{S}}_{u_{j}}\beta_{j}+\gamma^{\mathcal{T}}_{u_{j}}\alpha_{j})\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert$
Proof.
To simplify the derivation we first abuse the notation by letting $\mathbb{P}^{\mathcal{S}}$ denote $\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j,W_{u}=k)$ and $\mathbb{P}^{\mathcal{T}}$ denote $\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j,W_{u}=k)$.
$\lvert\gamma^{\mathcal{S}}_{u_{j}}\mathbb{P}^{\mathcal{S}}-\gamma^{\mathcal{T}}_{u_{j}}\mathbb{P}^{\mathcal{T}}\rvert-\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\cdot\mathbb{P}^{\mathcal{M}}$
$\leq\lvert\gamma^{\mathcal{S}}_{u_{j}}\mathbb{P}^{\mathcal{S}}-\gamma^{\mathcal{T}}_{u_{j}}\mathbb{P}^{\mathcal{T}}-(\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}})\mathbb{P}^{\mathcal{M}}\rvert$
$=\lvert\gamma^{\mathcal{S}}_{u_{j}}(\mathbb{P}^{\mathcal{S}}-\mathbb{P}^{\mathcal{M}})-\gamma^{\mathcal{T}}_{u_{j}}(\mathbb{P}^{\mathcal{T}}-\mathbb{P}^{\mathcal{M}})\rvert$
$\leq\gamma^{\mathcal{S}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{S}}-\mathbb{P}^{\mathcal{M}}\rvert+\gamma^{\mathcal{T}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{T}}-\mathbb{P}^{\mathcal{M}}\rvert$
In order to simplify the first term we substitute the definition of :-$\mathbb{P}^{\mathcal{M}}\coloneq\alpha_{j}\mathbb{P}^{\mathcal{S}}+\beta_{j}\mathbb{P}^{\mathcal{T}}$:
$\lvert\mathbb{P}^{\mathcal{S}}-\mathbb{P}^{\mathcal{M}}\rvert=\lvert\mathbb{P}^{\mathcal{S}}-\alpha_{j}\mathbb{P}^{\mathcal{S}}-\beta_{j}\mathbb{P}^{\mathcal{T}}\rvert=\lvert\beta_{j}\mathbb{P}^{\mathcal{S}}-\beta_{j}\mathbb{P}^{\mathcal{T}}\rvert$
Similarly for the second term:
$\lvert\mathbb{P}^{\mathcal{T}}-\mathbb{P}^{\mathcal{M}}\rvert=\lvert\mathbb{P}^{\mathcal{T}}-\alpha_{j}\mathbb{P}^{\mathcal{S}}-\beta_{j}\mathbb{P}^{\mathcal{T}}\rvert=\lvert\alpha_{j}\mathbb{P}^{\mathcal{T}}-\alpha_{j}\mathbb{P}^{\mathcal{S}}\rvert$
Using the two identities, we can proceed with the derivation:
$\lvert\gamma^{\mathcal{S}}_{u_{j}}\mathbb{P}^{\mathcal{S}}-\gamma^{\mathcal{T}}_{u_{j}}\mathbb{P}^{\mathcal{T}}\rvert-\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\cdot\mathbb{P}^{\mathcal{M}}$
$\leq\gamma^{\mathcal{S}}_{u_{j}}\beta_{j}\lvert\mathbb{P}^{\mathcal{S}}-\mathbb{P}^{\mathcal{T}}\rvert+\gamma^{\mathcal{T}}_{u_{j}}\alpha_{j}\lvert\mathbb{P}^{\mathcal{T}}-\mathbb{P}^{\mathcal{S}}\rvert$
$=(\gamma^{\mathcal{S}}_{u_{j}}\beta_{j}+\gamma^{\mathcal{T}}_{u_{j}}\alpha_{j})\lvert\mathbb{P}^{\mathcal{S}}-\mathbb{P}^{\mathcal{T}}\rvert$
See 3.3
Proof.
We start by converting the average error rate into individual error probabilities. For sufficiently large $|\mathcal{V}|$ we have $|\mathcal{V}|\approx|\mathcal{V}|-1$. Given $|\mathcal{V}^{\mathcal{S}}|\rightarrow\infty$ and $|\mathcal{V}^{\mathcal{T}}|\rightarrow\infty$, we can have $|\mathcal{V}|\approx|\mathcal{V}^{\mathcal{S}}|\approx|\mathcal{V}^{\mathcal{T}}|$. By applying the triangle inequality and assuming that the graphs are sufficiently large, we obtain the following inequality:
$\lvert\varepsilon^{\mathcal{S}}(h\circ g)-\varepsilon^{\mathcal{T}}(h\circ g)\rvert$
$=$ $\lvert\mathbb{P}^{\mathcal{S}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})-\mathbb{P}^{\mathcal{T}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})\rvert$
$=$ $\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}\neq Y_{u})-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}\neq Y_{u})\rvert$
To simplify the notation, define $\gamma^{\mathcal{U}}_{u_{j}}\coloneq\mathbb{P}^{\mathcal{U}}(Y_{u}=j)$, $\pi^{\mathcal{U}}_{v_{k}|u_{j}}\coloneq\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})$, and $\omega^{\mathcal{U}}_{u_{i}|u_{j},v_{k}}\coloneq\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ for $\mathcal{U})\in\left\{\mathcal{S},\mathcal{T}\right\}$. Using the the law of total probability and assuming that the label prediction depends on the distribution of neighborhood labels, we can derive the following identity:
$\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}\neq Y_{u})$ $=\sum_{i\neq j}\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i,Y_{u}=j)=\sum_{i\neq j}\mathbb{P}^{\mathcal{U}}(Y_{u}=j)\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j)$
$=\sum_{i\neq j}\mathbb{P}^{\mathcal{U}}(Y_{u}=j)\bigg{(}\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})\bigg{)}$
$=\sum_{i\neq j}\gamma^{\mathcal{U}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{U}}_{v_{k}|u_{j}}\omega^{\mathcal{U}}_{u_{i}|u_{j},v_{k}}$
The function of mean pooling naturally marginalizes the effect of degree magnitude, meaning that only the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself influence the prediction. We use Lemma A.1 to bound the terms above. Since $\forall j\in\mathcal{Y},\gamma^{\mathcal{S}}_{u_{j}}\text{ and }\gamma^{\mathcal{T}}_{u_{j}}\in[0,1]$ and we have $\alpha_{j}+\beta_{j}=1$, we obtain:
$|\varepsilon^{\mathcal{S}}(h\circ g)-\varepsilon^{\mathcal{T}}(h\circ g)|$
$\leq\sum_{i\neq j}\gamma^{\mathcal{S}}_{u_{j}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)$
$\leq\sum_{i\neq j}\lvert\gamma^{\mathcal{S}}_{u_{j}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\gamma^{\mathcal{T}}_{u_{j}}\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert$ (a)
$\leq\sum_{i\neq j}\left(\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\cdot\mathbb{P}^{\mathcal{M}}+(\gamma^{\mathcal{S}}_{u_{j}}\beta_{j}+\gamma^{\mathcal{T}}_{u_{j}}\alpha_{j})\lvert\mathbb{P}^{\mathcal{S}}-\mathbb{P}^{\mathcal{T}}\rvert\right)$ (b)
$=\sum_{i\neq j}\left(\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\cdot\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)+\gamma^{\mathcal{T}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert\right)$ (c)
$=\sum_{i\neq j}\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\cdot\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert$
$\leq\left(\sum_{j\in{\mathcal{Y}}}\lvert\gamma^{\mathcal{S}}_{u_{j}}-\gamma^{\mathcal{T}}_{u_{j}}\rvert\right)\cdot\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert$ (d)
$=2TV\left(\mathbb{P}^{\mathcal{S}}(Y_{u}=j),\mathbb{P}^{\mathcal{T}}(Y_{u}=j)\right)\cdot\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert$ (e)
By applying the triangle inequality, (a) is obtained. Following Lemma A.1 we have (b). (c) is obtained by choosing $\alpha_{j}=1$ and $\beta_{j}=0,\forall j\in\mathcal{Y}$. (d) is derived by applying Holder’s inequality. (e) is based on the definition of total variation distance where $TV(\mathbb{P}^{\mathcal{S}}(Y_{u}=j),\mathbb{P}^{\mathcal{T}}(Y_{u}=j))=\frac{1}{2}\sum_{j\in{\mathcal{Y}}}\lvert\mathbb{P}^{\mathcal{S}}(Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(Y_{u}=j)\rvert$ is the total variation distance between the label distribution of source and target.
Similarly, the second term can be decomposed by Lemma A.1 while setting $\omega^{\mathcal{M}}_{u_{i}|u_{j},v_{k}}\coloneq\alpha_{j}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})+\beta_{j}\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$, $\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}\coloneq\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$, and $\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\coloneq\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$. The decomposition is instead with respect to $\pi^{\mathcal{U}}_{v_{k}|u_{j}}\coloneq\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})$.
$\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\lvert\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j)\rvert$
$=\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\begin{aligned} &\lvert\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j,\{Y_{v}:v\in\mathcal{N}_{u}\})\\ &\quad-\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j,\{Y_{v}:v\in\mathcal{N}_{u}\})\rvert\end{aligned}$
$\leq\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\lvert\pi^{\mathcal{S}}_{v_{k}|u_{j}}\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\pi^{\mathcal{T}}_{v_{k}|u_{j}}\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\rvert$
$\leq\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\left(\lvert\pi^{\mathcal{S}}_{v_{k}|u_{j}}-\pi^{\mathcal{T}}_{v_{k}|u_{j}}\rvert\cdot\omega^{\mathcal{M}}_{u_{i}|u_{j},v_{k}}+(\pi^{\mathcal{S}}_{v_{k}|u_{j}}\beta_{j}+\pi^{\mathcal{T}}_{v_{k}|u_{j}}\alpha_{j})\lvert\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\rvert\right)$ (a)
$=\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\lvert\pi^{\mathcal{S}}_{v_{k}|u_{j}}-\pi^{\mathcal{T}}_{v_{k}|u_{j}}\rvert\cdot\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{T}}_{v_{k}|u_{j}}\ \lvert\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\rvert$ (b)
$=\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{S}}_{v_{k}|u_{j}}\lvert 1-\frac{\pi^{\mathcal{T}}_{v_{k}|u_{j}}}{\pi^{\mathcal{S}}_{v_{k}|u_{j}}}\rvert\cdot\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{T}}_{v_{k}|u_{j}}\ \lvert\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\rvert$
$\leq\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\left(\max\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}\rvert\right)\cdot\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{T}}_{v_{k}|u_{j}}\ \lvert\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\rvert$ (c)
$\leq\mathbb{E}^{\mathcal{T}}\left[\left(\max\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}\rvert\right)\right]\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})+\sum_{i\neq j}\gamma^{\mathcal{T}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{T}}_{v_{k}|u_{j}}\ \lvert\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}\rvert$
$\leq\mathbb{E}_{Y_{u}}^{\mathcal{T}}\left[\left(\max\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}\rvert\right)\right]\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})+\max_{i\ \neq j}\mathbb{E}^{\mathcal{T}}\left[|\omega^{\mathcal{S}}_{u_{i}|u_{j},v_{k}}-\omega^{\mathcal{T}}_{u_{i}|u_{j},v_{k}}|\right]$
$\leq\mathbb{E}_{Y_{u}}^{\mathcal{T}}&\left[\left(\max\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}\rvert\right)\right]\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\\ &+\max_{i\neq j}\mathbb{E}^{\mathcal{T}}_{Y_{v}}[|\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})-\mathbb{P}^{\mathcal{T}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})|]\end{aligned}$
$\leq\mathbb{E}_{Y_{u}}^{\mathcal{T}}\left[\left(\max\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})}\rvert\right)\right]\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})+\Delta_{CE}$ (d)
(a) comes from Lemma A.1. (b) comes from setting $\alpha_{j}=1$ and $\beta_{j}=0,\forall j\in\mathcal{Y}$ (c) The inequality comes from Holder’s inequality, where we take the maximum-norm for $k$ and take the $L_{1}$ norm to marginalize $Y_{v}$: $\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)=\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ (d) The last term in the previous line corresponds to the conditional error gap in (Tachet des Combes et al., 2020) but in a GNN mean pooling fashion.
The term $\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ can be interpreted as the probability of error prediction using a mean pooling GNN encoder. By definition, $Y_{v}$ represents a random variable of the neighborhood label ratio and $Y_{u}$ denotes a random variable of self class label, their realization corresponds to the mean pooled and center value. As a reult, $\Delta_{CE}$ only exists if and only if feature shift exists. Putting everything together, we have
$\lvert\varepsilon^{\mathcal{S}}(g\circ\phi)-\varepsilon^{\mathcal{T}}(g\circ\phi)\rvert$
$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Neighborhood shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature shift}}\vspace{-2mm}$
Appendix B Additional Results
B.1 Results of GPRGNN
Table 5 presents the results of the GPRGNN backbone on the MAG dataset. Compared to GraphSAGE, we observe that GPRGNN is less robust under structure shifts due to its linear combination of k-hop representations. We also notice a similar tendency as with GraphSAGE, where T3A outperforms all other baselines. TSA slightly improves upon T3A by refining the decision boundary. However, we frequently observe a performance decrease when combining AdaRC with T3A. This drop is attributed to the fact that their proposed PIC loss relies on minimizing intra-variance, which depends on the quality of node hidden representations on the target domain. Under imbalanced training and label shift, node hidden representations can easily degrade, causing AdaRC performance to become unstable.
Table 5: MAG GPRGNN results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
[…]
ERM 26.26$\pm$1.53 27.09$\pm$0.74 38.90$\pm$1.06 22.88$\pm$1.14 20.69$\pm$0.64 27.58$\pm$0.55 10.77$\pm$0.55 17.41$\pm$0.86 12.49$\pm$0.74 9.05$\pm$0.42
GTrans 26.04$\pm$1.41 26.96$\pm$0.72 38.71$\pm$0.96 23.07$\pm$1.08 20.80$\pm$0.70 27.49$\pm$0.46 10.97$\pm$0.56 17.41$\pm$0.78 12.58$\pm$0.75 9.08$\pm$0.41
SOGA 34.70$\pm$0.71 35.18$\pm$0.33 48.47$\pm$0.30 33.63$\pm$0.73 27.98$\pm$0.42 35.16$\pm$0.57 18.50$\pm$0.45 28.35$\pm$1.04 20.11$\pm$0.71 14.73$\pm$0.58
TENT 13.79$\pm$0.72 25.67$\pm$0.50 29.13$\pm$0.80 18.00$\pm$0.14 20.85$\pm$0.26 16.29$\pm$0.82 7.59$\pm$0.73 13.81$\pm$0.74 7.82$\pm$0.74 5.94$\pm$0.34
LAME 24.54$\pm$2.55 22.50$\pm$0.85 35.97$\pm$1.42 17.61$\pm$1.46 15.71$\pm$0.92 28.02$\pm$0.55 9.08$\pm$0.61 15.58$\pm$1.02 10.87$\pm$0.70 8.01$\pm$0.38
T3A 40.58$\pm$0.23 39.00$\pm$0.85 48.96$\pm$1.00 46.13$\pm$1.56 32.35$\pm$1.29 38.92$\pm$2.23 38.07$\pm$1.45 46.23$\pm$1.73 42.82$\pm$1.52 28.65$\pm$0.65
AdaRC-TENT 13.90$\pm$0.71 25.87$\pm$0.47 29.30$\pm$0.82 18.14$\pm$0.19 21.03$\pm$0.26 16.38$\pm$0.80 7.62$\pm$0.73 13.84$\pm$0.74 7.85$\pm$0.81 5.97$\pm$0.36
AdaRC-LAME 8.93$\pm$4.76 8.18$\pm$0.78 15.46$\pm$1.34 6.25$\pm$1.43 6.69$\pm$0.81 1.35$\pm$2.69 17.60$\pm$8.89 33.42$\pm$2.41 18.81$\pm$1.42 4.23$\pm$8.02
AdaRC-T3A 22.34$\pm$18.21 32.18$\pm$1.70 42.85$\pm$1.15 34.77$\pm$2.62 25.75$\pm$0.62 45.27$\pm$2.57 33.01$\pm$1.17 42.08$\pm$0.82 38.57$\pm$1.50 29.30$\pm$0.47
TSA-TENT 14.25$\pm$0.68 25.76$\pm$0.51 29.30$\pm$0.79 18.30$\pm$0.26 21.20$\pm$0.20 16.57$\pm$0.78 7.69$\pm$0.72 14.03$\pm$0.70 8.07$\pm$0.79 6.03$\pm$0.39
TSA-LAME 26.33$\pm$2.25 22.76$\pm$1.44 35.57$\pm$1.43 18.15$\pm$2.08 16.94$\pm$1.06 29.70$\pm$0.45 11.04$\pm$0.61 17.90$\pm$0.89 12.46$\pm$0.43 9.32$\pm$0.51
TSA-T3A 40.81$\pm$0.23 39.13$\pm$0.77 49.16$\pm$0.85 47.05$\pm$1.74 33.27$\pm$1.21 41.06$\pm$3.34 38.34$\pm$1.25 46.38$\pm$1.83 43.64$\pm$0.85 28.71$\pm$0.57
B.2 Results of Ablation Study
Table 6: Ablation Study on TSA.
[…]
TENT+TSA Base 27.30$\pm$1.61 32.84$\pm$0.78 40.82$\pm$0.99 32.53$\pm$0.91 28.62$\pm$0.63 27.89$\pm$0.98 16.22$\pm$1.17 24.87$\pm$0.56 22.05$\pm$0.84 14.20$\pm$0.71
w/o Nbr. Align 27.18$\pm$1.70 32.75$\pm$0.69 40.82$\pm$0.99 32.26$\pm$0.95 28.41$\pm$0.59 27.53$\pm$0.92 16.22$\pm$1.17 24.87$\pm$0.56 22.05$\pm$0.84 13.95$\pm$0.76
w/o SNR 27.00$\pm$1.21 32.76$\pm$0.75 40.80$\pm$0.91 32.52$\pm$0.90 28.57$\pm$0.59 27.70$\pm$0.88 15.84$\pm$0.85 24.62$\pm$0.48 21.37$\pm$0.73 14.11$\pm$0.60
LAME+TSA Base 37.95$\pm$0.97 36.29$\pm$1.65 46.86$\pm$1.13 32.86$\pm$2.23 27.22$\pm$1.48 44.83$\pm$0.88 28.51$\pm$0.44 39.80$\pm$0.99 24.54$\pm$0.87 22.39$\pm$0.30
w/o Nbr. Align 37.90$\pm$0.97 36.08$\pm$1.73 46.75$\pm$0.88 32.70$\pm$2.29 26.49$\pm$1.76 44.83$\pm$0.88 28.41$\pm$0.36 39.72$\pm$0.99 24.24$\pm$0.75 22.23$\pm$0.32
w/o SNR 35.35$\pm$0.84 33.45$\pm$1.60 44.51$\pm$1.22 29.89$\pm$1.94 24.41$\pm$1.43 40.17$\pm$1.11 22.44$\pm$0.99 32.57$\pm$1.46 17.33$\pm$0.47 17.07$\pm$0.61
T3A+TSA Base 41.65$\pm$0.99 47.01$\pm$2.08 51.65$\pm$0.90 46.61$\pm$0.88 43.45$\pm$0.81 48.09$\pm$0.60 39.18$\pm$1.87 46.50$\pm$0.25 43.70$\pm$1.38 30.89$\pm$2.13
w/o Nbr. Align 41.64$\pm$1.10 46.37$\pm$2.11 50.32$\pm$1.11 46.44$\pm$0.80 42.09$\pm$1.26 46.47$\pm$1.27 39.12$\pm$1.85 46.23$\pm$0.61 43.60$\pm$1.21 30.92$\pm$2.03
w/o SNR 41.59$\pm$0.94 45.54$\pm$1.81 51.55$\pm$0.95 46.51$\pm$0.94 41.35$\pm$1.09 48.11$\pm$0.62 38.71$\pm$1.00 46.29$\pm$0.26 43.12$\pm$0.76 29.90$\pm$1.35
B.3 Visualization of Distribution Shift
(a) Source (b) Feature Shift (c) Neighborhood Shift (d) Label Shift (e) SNR
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: The t-SNE visualization of a one-layer GSN under different distribution shifts: (top) the mean-pooled neighborhood hidden representations and (bottom) the node representations before passing through the classifier, positioned along its decision boundary. (a) indicates the source domain where the model is pretrained. (b), (c), and (d) stand for feature shift, neighborhood shift, and label shift. (e) represent the impact of SNR induced by degree shift. The color of the nodes represents the ground-truth labels. The legends in the bottom plots show the accuracy and recall scores for each class. The red contours in the bottom (d) highlight the decision boundary of the minority class.
Appendix C Dataset Details
C.1 Dataset Statistics
Here we summarize the statistics of the four real-world graph datasets. The edges are counted twice in the edge list as they are undirected graphs.
Table 7: Arxiv and ACM/DBLP dataset statistics
ACM DBLP 1950-2007 1950-2009 1950-2011 1950-2016 1950-2018
Nodes $7410$ $5578$ $4980$ $9410$ $17401$ $69499$ $120740$
Edges $11135$ $7341$ $5849$ $13179$ $30486$ $232419$ $615415$
Features $7537$ $7537$ $128$ $128$ $128$ $128$ $128$
Classes $6$ $6$ $40$ $40$ $40$ $40$ $40$
Table 8: MAG dataset statistics
US CN DE JP RU FR
Nodes $132558$ $101952$ $43032$ $37498$ $32833$ $29262$
Edges $697450$ $285561$ $126683$ $90944$ $67994$ $78222$
Features $128$ $128$ $128$ $128$ $128$ $128$
Classes $20$ $20$ $20$ $20$ $20$ $20$
Table 9: Pileup dataset statistics
gg-10 qq-10 gg-30 qq-30 gg-50 gg-140
Nodes $18611$ $17242$ $41390$ $38929$ $60054$ $154750$
Edges $53725$ $42769$ $173392$ $150026$ $341930$ $2081229$
Features $28$ $28$ $28$ $28$ $28$ $28$
Classes $2$ $2$ $2$ $2$ $2$ $2$
C.2 Dataset Shift Statistics
Below we provide two metrics indicating the neighborhood and label shift in real-world dataset. We measure the neighborhood shift by computing the weighted average of the total variation distance between of the neighborhood distribution:
[…]
The label shift is measured as the total variation distance between the label distribution, which corresponds to the first term in Theorem 3.3.
[…]
Table 10: MAG dataset shift metrics
$US\rightarrow CN$ $US\rightarrow DE$ $US\rightarrow JP$ $US\rightarrow RU$ $US\rightarrow FR$ $CN\rightarrow US$ $CN\rightarrow DE$ $CN\rightarrow JP$ $CN\rightarrow RU$ $CN\rightarrow FR$ Nbr. Shift $0.2137$ $0.1902$ $0.1476$ $0.2855$ $0.2238$ $0.1946$ $0.2053$ $0.1361$ $0.2025$ $0.2039$ Label Shift $0.2734$ $0.1498$ $0.1699$ $0.3856$ $0.1706$ $0.2734$ $0.2691$ $0.1522$ $0.2453$ $0.2256$
Table 11: Pileup dataset shift metrics
Pileup Conditions Physical Processes Domains $\text{PU}10\rightarrow 30$ $\text{PU}30\rightarrow 10$ PU1050105010$\rightarrow$ 5010 → 50 PU5010501050$\rightarrow$ 1050 → 10 PU301403014030$\rightarrow$ 14030 → 140 PU1403014030140$\rightarrow$ 30140 → 30 $gg\rightarrow qq$ $qq\rightarrow gg$ Nbr. Shift $0.1493$ $0.1709$ $0.2039$ $0.2595$ $0.1293$ $0.1635$ $0.0214$ $0.0205$ Label Shift $0.2425$ $0.2425$ $0.2982$ $0.2982$ $0.1377$ $0.1377$ $0.0341$ $0.0348$
Table 12: Real dataset shift metrics
1950-2007 1950-2009 1950-2011 DBLP/ACM Domains $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $D\rightarrow A$ $A\rightarrow D$ Nbr. Shift $0.3005$ $0.3450$ $0.2583$ $0.3120$ $0.1833$ $0.2567$ $0.1573$ $0.2227$ Label Shift $0.2938$ $0.4396$ $0.2990$ $0.4552$ $0.2853$ $0.4438$ $0.3434$ $0.3435$
Appendix D Detailed Experimental Setup
D.1 Datasets Setup
For Arxiv, MAG, and DBLP/ACM we follow the setting of (Liu et al., 2024). Hence we discuss the setting of synthetic dataset and pileup in the following:
Synthetic Dataset.
The synthetic dataset is generated by the contextual stochastic block model (CSBM) (Deshpande et al., 2018). The generated graph contains 6000 nodes and 3 classes. We alter the edge connection probability matrix $\mathbf{B}$ and the label ratio $\mathbb{P}_{Y}$ Specifically, the edge connection probability matrix $\mathbf{B}$ is a symmetric matrix given by
$\mathbf{B}=\begin{bmatrix}p&q&q\\ q&p&q\\ q&q&p\end{bmatrix}$
where $p$ and $q$ indicate the intra- and the inter-class edge probability respectively. We assume the node features are sampled from a Gaussian distribution $X_{u}\sim{\mathcal{N}}(\bm{\mu}_{u},\sigma{\bm{I}})$ where we set $\sigma=0.3$ as well as $\bm{\mu}_{u}=[1,0,0]$ for class 0, $\bm{\mu}_{u}=[0,1,0]$ for class 1, and $\bm{\mu}_{u}=[0,0,1]$ for class 2.
  • The source graph for setting 1-6 has $\mathbb{P}_{Y}=\left[0.1,0.3,0.6\right]$, with $p=0.01$ and $q=0.0025$.
  • For neighborhood shift we fix the same class ratio but change the connection probability:
    • Condition 1: $p=0.005$, $q=0.00375$.
    • Condition 2: $p=0.005$, $q=0.005$.
  • For condition 3 and 4 we introduce degree shift for SNR discrepancy.
    • Condition 3: $p=\frac{0.005}{2}$, $q=\frac{0.00375}{2}$.
    • Condition 4: $p=\frac{0.005}{2}$, $q=\frac{0.005}{2}$.
  • For condition 5 and 6 we investigate structure shift under training from give training is from the imbalanced source label ratio.
    • Condition 5: $\mathbb{P}_{Y}=\left[1/3,1/3,1/3\right]$ $p=\frac{0.005}{2}$, $q=\frac{0.00375}{2}$.
    • Condition 6: $\mathbb{P}_{Y}=\left[1/3,1/3,1/3\right]$ $p=\frac{0.005}{2}$, $q=\frac{0.005}{2}$.
  • The source graph for Condition 7-8 has $\mathbb{P}_{Y}=\left[1/3,1/3,1/3\right]$, with $p=0.01$ and $q=0.0025$.
  • For condition 7 and 8 we investigate structure shift under training from balanced source ratio. The $p$ and $q$ in Condition 7 and 8 is the same as in Condition 5 and 6, respectively.
    • Condition 7: $\mathbb{P}_{Y}=[0.1,0.3,0.6]$, $p=\frac{0.005}{2}$, $q=\frac{0.00375}{2}$.
    • Condition 8: $\mathbb{P}_{Y}=[0.1,0.3,0.6]$, $p=\frac{0.005}{2}$, $q=\frac{0.005}{2}$.
Pileup.
Under the study of different PU levels we use the data from signal $gg$. Compared to the Pileup datasets used in the (Liu et al., 2024), we explicitly label all types of particles to better align the neighborhood distribution. Note that the charged particle can achieve all most perfect recall as their label information is encoded in their feature, hence the four-class classification still reduces into a binary classification between LC neutral particles and OC neutral particles. For the study of different data generation process, we use PU10 from signal $gg\rightarrow qq$ and $qq\rightarrow gg$.
D.2 Pretraining Setup
Model architecture and Pretraining
We use GraphSAGE as the model backbone, with a 3-layer mean pooling GNN as the encoder and a 2-layer MLP with a batch normalization layer as the classifier. The hidden dimension of the encoder is set to 300 for Arxiv and MAG, 50 for Pileup, 128 for DBLP/ACM, and 20 for the synthetic datasets. The hidden dimension of the classifier is set to 300 for Arxiv and MAG, 50 for Pileup, 40 for DBLP/ACM, and 20 for the synthetic datasets. For the GPRGNN backbone, we follow the same configuration but increase the number of encoder layers to 5. All models are trained for 400 epochs with the learning rate set to 0.003. A learning rate scheduler is applied with a decay rate of 0.9. All experiments are repeated 5 times under different initializations and data splits to ensure consistency.
D.3 Evaluation and Metric
The source graph is splitted into train/val/test 60/20/20 during the pretraining stage. The target graph is splitted into labeled/unlabeled 3/97 during test time. We use the the labeled nodes in the target graph to do hyperparameter tuning and select the model with the optimal validation score. We follow the metrics in (Liu et al., 2024) for evaluation. MAG, Arxiv, DBLP/ACM, and synthetic datasets are evaluated with accuracy. For the MAG datasets, only the top 19 classes are evaluated. Pileup is evaluated with f1 score.
D.4 Hyperparameter Tuning
Below we introduce the search space of the hyperparameters. LAME is a parameter-free approach so we do not tune it.
  • TSA introduces three hyperparameters (1) the learning rate $lr$ for optimizing $\alpha$, (2) the ratio $\rho_{1}$ for reliable assignment of $\gamma$ based on entropy $H(\hat{y})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$, and (3) the ratio $\rho_{2}$ for filtering out unreliable hard pseudo-labels in Eq. 4 based on entropy $H(\hat{y})\leq\rho_{2}\cdot\ln(|{\mathcal{Y}}|)$. We select $lr$ from $\{0.001,0.01,0.05,0.1\}$, $\rho_{1}$ from $\{0,0.01,0.1,0.5\}$, and $\rho_{2}$ from $\{0.1,1.0\}$. We present the results updated by one epoch. We observe that given the same domain shifts, the optimal hyperparameters of $\rho_{1}$ may differ due to different performance of the boundary refinement approaches (TENT, T3A, and LAME).
  • AdaRC introduces two hyperparameters learning rate $lr$ and the number of epochs $T$. Based on their hyperparameter study, We select $lr$ from $\{1,5,10\}$ and $T$ from $\{1,10,100\}$.
  • GTrans introduces four hyperparameters learning rate of feature adaptation $lr_{f}$, learning rate of structure adaptation $lr_{A}$, the number of epochs $T$, and the budget to update. We select $lr_{f}$ from $\{0.001,0.01\}$, $lr_{A}=\{0.01,0.1,0.5\}$ from $\{0.01,0.05\}$, $T$ from $\{5,10\}$, and budget from $\{0.01,0.05,0.1\}$.
  • SOGA introduces one hyperparameters $lr$ for one epoch update. We select $lr$ from $\{0.001,0.01\}$.
  • TENT introduces one hyperparameters $lr$ for one epoch update. We select $lr$ from $\{0.001,0.01,0.05\}$.
  • T3A introduces one hyperparameters $M$ for deciding the number of supports to restore. We select $M$ from $\{5,20,50,100\}$

### v2

@@
 Abstract
-Graph-based learning has achieved remarkable success in domains ranging from recommendation to fraud detection and particle physics by effectively capturing underlying interaction patterns.
-However, it often struggles to generalize when distribution shifts occur, particularly those involving changes in network connectivity or interaction patterns.
-Existing approaches designed to mitigate such shifts typically require retraining with full access to source data, rendering them infeasible under strict computational or privacy constraints.
-To address this limitation, we propose a test-time structural alignment (TSA) algorithm for Graph Test-Time Adaptation (GTTA), a novel method that aligns graph structures during inference without revisiting the source domain.
-Built upon a theoretically grounded treatment of graph data distribution shifts, TSA integrates three key strategies: an uncertainty-aware neighborhood weighting that accommodates structure shifts, an adaptive balancing of self-node and neighborhood-aggregated representations driven by node representations’ signal-to-noise ratio, and a decision boundary refinement that corrects remaining label and feature shifts.
-Extensive experiments on synthetic and real-world datasets demonstrate that TSA can consistently outperform both non-graph TTA methods and state-of-the-art GTTA baselines.
-Machine Learning, ICML
-Figure 1: Example of the distribution shifts of neighborhood information due to different experimental conditions in the LHC experiments.
-The goal of pileup mitigation is to detect LC neutral particles.
-The particles are modeled by kNN graphs (dashed circles in the figure) to use nearby particles for drawing inference.
-The model is trained on ${\mathcal{G}}^{\mathcal{S}}$ but needs to generalize to ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$ over time.
-The inferred nodes within the circles are the LC neutral particles, but their neighborhood node label ratio changes in ${\mathcal{G}}^{{\mathcal{T}}_{1}}$, ${\mathcal{G}}^{{\mathcal{T}}_{2}}$.
-In ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ the homophily ratio changes as one of the neighbor node is an LC neutral particle.
-Both cases represent neighborhood shift which this work aims to address.
-A more formal definition of these shifts is discussed in Sec.
-3.
+Graph-based learning excels at capturing interaction patterns in diverse domains like recommendation, fraud detection, and particle physics.
+However, its performance often degrades under distribution shifts, especially those altering network connectivity.
+Current methods to address these shifts typically require retraining with the source dataset, which is often infeasible due to computational or privacy limitations.
+We introduce Test-Time Structural Alignment (TSA), a novel algorithm for Graph Test-Time Adaptation (GTTA) that aligns graph structures during inference without accessing the source data.
+Grounded in a theoretical understanding of graph data distribution shifts, TSA employs three synergistic strategies: uncertainty-aware neighborhood weighting to accommodate neighbor label distribution shifts, adaptive balancing of self-node and aggregated neighborhood representations based on their signal-to-noise ratio, and decision boundary refinement to correct residual label and feature shifts.
+Extensive experiments on synthetic and real-world datasets demonstrate TSA’s consistent outperformance of both non-graph TTA methods and state-of-the-art GTTA baselines.
 1 Introduction
-Graph-based methods have become indispensable in handling structured data across a wide range of real-world applications (Duvenaud et al., 2015; Bronstein et al., 2017; Zhang et al., 2019a; Stokes et al., 2020), achieving significant success when training and testing data originate from similar distributions.
+Graph-based methods have become indispensable in handling structured data across a wide range of real-world applications [1, 2, 3, 4], achieving significant success when training and testing data originate from similar distributions.
 However, these methods struggle to generalize to test-time data in a different domain, where variations in time, location, or experimental conditions result in distinct graph connection patterns.
-Although some literature on graph domain adaptation (GDA) (Wu et al., 2020; You et al., 2023; Zhu et al., 2021; Liu et al., 2023) seeks to bridge these gaps by aligning labeled source distributions with target distributions, such approaches are often infeasible in practice due to the cost of retraining and the limited availability of source data in the test time.
-Figure 2: TSA utilizes neighborhood alignment and SNR-inspired adjustment to mitigate neighborhood information discrepancy.
-It further adjusts the decision boundary to get refined predictions $\hat{y}$.
-The refined soft pseudo-labels are used to estimate parameter $\bm{\gamma}$ for neighborhood alignment and to optimize $\alpha$ for combining self representations with neighborhood aggregated representations.
-Real-world applications may operate under constraints such as lightweight computation, limited storage, and strict privacy requirements, which make reprocessing of source data generally infeasible (Wang et al., 2021c; Wu et al., 2021).
-For example, in particle tracking based on graph learning at the Large Hadron Collider (LHC) (Shlomi et al., 2020; Highfield, 2008; Miao et al., 2024), the connections between particles differ as experimental conditions change over time.
-The tracker is expected to be capable of adapting on-the-fly (Li et al., 2023; Komiske et al., 2017).
-Similarly, in fraud detection for financial networks  (Clements et al., 2020; Wang et al., 2021b), privacy-sensitive training data may not be accessed during testing, while transaction patterns may change between different regions  (Wang et al., 2019; Dou et al., 2020).
-These scenarios indicate the need for graph test-time domain adaptation (GTTA), which enables models to adjust to new test domains without re-accessing the source domain or incurring the overhead of full retraining.
+Although some literature on graph domain adaptation (GDA) [5, 6, 7, 8] seeks to bridge these gaps by aligning labeled source distributions with target distributions, such approaches are often infeasible in practice due to the cost of retraining and the limited availability of source data at the test time.
+Real-world applications may operate under constraints such as lightweight computation, limited storage, and strict privacy requirements, which make reprocessing of source data generally infeasible [9, 10].
+For example, Fig.
+1 illustrates pileup mitigation based on graph learning at the Large Hadron Collider (LHC) [11, 12, 13], where the connections between particles differ as experimental conditions change.
+To address this, the detector is expected to be capable of adapting on-the-fly [14, 15], potentially on embedded hardware.
+Similarly, in fraud detection for financial networks [16, 17], privacy-sensitive training data may not be accessed during testing, while transaction patterns may change between different regions [18, 19].
+These scenarios indicate the need for graph test-time domain adaptation (GTTA), which enables models to adjust to new test domains without re-accessing the massive source dataset during adaptation or incurring the overhead of full retraining.
 Accurate node classification in graph-structured data relies heavily on effectively leveraging neighborhood information.
-Graph neural networks (GNNs) (Kipf & Welling, 2017; Veličković et al., 2018; Hamilton et al., 2017; Hamilton, 2020) have achieved significant success across a wide range of applications by utilizing this neighborhood context, yet they remain sensitive to distribution shifts in node neighborhoods (Gui et al., 2022; Ji et al., 2023).
-As shown in Fig.
-1, shifts in LHC data can lead to distinct neighborhood connections among center particles in the same class (neighborhood shift).
-Existing methods for test-time adaptation (TTA) often focus on adjusting only the classifier’s final layer (Iwasawa & Matsuo, 2021), modifying its outputs (Boudiaf et al., 2022), or adapting normalization statistics (Wang et al., 2021a; Niu et al., 2023), primarily targeting image-based tasks with independently labeled data points.
+Graph neural networks (GNNs) [20, 21, 22, 23] have achieved significant success across a wide range of applications by utilizing this neighborhood context, yet they remain sensitive to distribution shifts in node neighborhoods [24, 25].
+In the above LHC example illustrated by Fig.
+1, distribution shifts lead to distinct neighborhood connections among center particles in the same class, causing a neighborhood shift.
+Existing methods for test-time adaptation (TTA) often focus on adjusting only the classifier’s final layer [26], modifying its outputs [27], or adapting normalization statistics [28, 29], primarily targeting image-based tasks with independently labeled data points.
 Consequently, they may not adequately address neighborhood shifts for node classification in graphs.
-While recent works on GTTA attempt to address the structural changes using graph augmentation (Jin et al., 2023), self-supervised learning (Mao et al., 2024).
-These approaches mainly rely on heuristics and homophily consistency, failing to fully address changes in neighboring node connections in principle.
+While recent works on GTTA attempt to address the structural changes using graph augmentation [30], self-supervised learning [31, 32], these approaches mainly rely on heuristics and homophily consistency, failing to fully address changes in neighboring node connections in principle.
 In this work, we propose a model-agnostic framework with test-time structural alignment, named TSA, as illustrated in Fig.
@@
 We begin by examining the generalization gap in GTTA for a pretrained source model.
-Our theoretical findings, further validated by empirical studies on synthetic datasets, indicate the importance of aligning the label distributions of neighboring nodes among the same class of nodes across the source and target domains.
-In the context of training-time GDA, PairAlign (Liu et al., 2024) employs a neighborhood weighting strategy to recalibrate the influence of neighboring nodes during message aggregation for a similar alignment.
-However, extending this approach to GTTA necessitates our more nuanced investigation, leading to the following contributions:
-Firstly, the weight assignment relies on knowledge of node labels, which are unavailable in the target graph under GTTA.
+Our theoretical findings, further validated by empirical studies, indicate the importance of aligning the neighborhood distributions across the source and target domains.
+Our contributions are summarized as follows:
+Firstly, TSA employs a neighborhood weighting strategy in GTTA to recalibrate the influence of neighboring nodes during message aggregation, thereby enabling structural alignment under the GTTA setting.
+However, the weight assignment relies on knowledge of node labels, which are unavailable in the target graph.
 Assigning weights based on pseudo labels may, in fact, degrade performance.
@@
 Additionally, TSA optimizes the test-time combination of self-node representations and neighborhood-aggregated representations based on their signal-to-noise ratio (SNR).
-For instance, when the target graph is sparse, TSA reduces reliance on neighborhood messages due to their high variance and low SNR.
-Conversely, as layer depth increases, it assigns greater weight to self-node representations, which become progressively denoised.
-Lastly, TSA integrates non-graph TTA methods to refine the decision boundary, mitigating mismatches caused by additional label and feature shifts once neighborhood shift has been addressed.
+Our experiments show that neighborhood-aggregated representations tend to have higher SNR and are better denoised when the neighbor set size and the GNN layer depth increase.
+Under both conditions, TSA adjusts the combination by assigning greater weight to neighborhood-aggregated representations than to self-node representations to enhance model performance.
+Lastly, TSA integrates non-graph TTA methods to refine the decision boundary, mitigating mismatches caused by additional label and feature shifts once neighborhood shifts have been addressed.
 We conduct extensive experiments on synthetic and four real-world datasets, including those from high-energy physics and citation networks, with multiple GNN backbones, to demonstrate the effectiveness of TSA.
-TSA outperforms non-graph TTA baselines by up to 12% on synthetic datasets and surpasses all non-graph baselines.
-Compared to existing GTTA baselines, TSA achieves an average of 10% improvement on all real-world datasets.
+TSA outperforms non-graph TTA baselines by up to 21.25% on synthetic datasets and surpasses all non-graph baselines.
+Compared to existing GTTA baselines, TSA achieves an average improvement of 10% on all real-world datasets.
+Figure 1: Example of the distribution shifts of neighborhood information due to different experimental conditions.
+Pileup mitigation aims to detect LC neutral particles using kNN graphs (dashed circles) that leverage nearby particles for inference.
+The model is trained on ${\mathcal{G}}^{\mathcal{S}}$ but needs to generalize to ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$.
+The inferred nodes within the circles are the LC neutral particles, but their neighborhood node label ratios change in ${\mathcal{G}}^{{\mathcal{T}}_{1}}$ and ${\mathcal{G}}^{{\mathcal{T}}_{2}}$.
+Both cases represent neighborhood shift, which this work aims to address.
+A more formal definition of these shifts is discussed in Sec.
+3.
+Figure 2: TSA utilizes neighborhood alignment to address neighborhood shift and SNR-inspired adjustment to mitigate SNR shift.
+It further adjusts the decision boundary to get refined predictions $\hat{y}$.
+The refined soft pseudo-labels are used to estimate the parameter $\bm{\gamma}$ for neighborhood alignment and to optimize $\alpha$ for combining self representations with neighborhood-aggregated representations.
 2 Preliminaries and Related Works
 2.1 Notations and Problem Setup
-We use upper-case letters, such as $Y$ to represent random variables, lower-case letters, such as $y$ to represent their realization.
-The calligraphic letters, such as ${\mathcal{Y}}$ denote the domain of random variables.
-We use bold capital such as ${\mathbf{Y}}$ to represent the vectorized corresponds, i.e., a collections of random variables.
-The probability distribution of a random variable $Y$ for a realization is expressed as $\mathbb{P}(Y=y)$.
+We use upper-case letters, such as $Y$ to represent random variables, and lower-case letters, such as $y$ to represent their realizations.
+Calligraphic letters, such as ${\mathcal{Y}}$ denote the domain of random variables.
+We use bold capital letters such as ${\mathbf{Y}}$ to represent the vectorized counterparts, i.e., collections of random variables.
+The probability of a random variable $Y$ for a realization is expressed as $\mathbb{P}(Y=y)$.
 Graph Neural Networks (GNNs).
@@
 GNNs utilize neighborhood information by encoding ${\bm{A}}$ and ${\bm{X}}$ into node representations $\{h_{v}^{(k)},v\in\mathcal{N}_{u}\}$.
-With $h_{u}^{(1)}=x_{u}$, the message passing in standard GNNs for node $v$ and each layer $k\in[L]\coloneqq\{1,\dots,L\}$ can be written as
+With $h_{u}^{(1)}=x_{u}$, the message passing in standard GNNs for node $u$ and each layer $k\in[L]\coloneqq\{1,\dots,L\}$ can be written as
 $h_{u}^{(k+1)}=\text{UPT}(h_{u}^{(k)},\text{AGG}(\{h_{v}^{(k)},v\in\mathcal{N}_{u}\}))$ (1)
-where $\mathcal{N}_{u}$ denotes the set of neighbors of node $u$, which $|\mathcal{N}_{u}|$ represent the node degree $d_{u}$.
+where $\mathcal{N}_{u}$ denotes the set of neighbors of node $u$, and $|\mathcal{N}_{u}|$ represents the node degree $d_{u}$.
 The AGG function aggregates messages from the neighbors, and the UPT function updates the node representations.
+Test-Time Adaptation (TTA).
+Assume the model consists of a feature encoder $\phi:{\mathcal{X}}\rightarrow{\mathcal{H}}$ and a classifier $g:{\mathcal{H}}\rightarrow{\mathcal{Y}}$ and each domain ${\mathcal{U}}\in\{{\mathcal{S}},{\mathcal{T}}\}$ consists of a joint feature and label distribution $\mathbb{P}^{\mathcal{U}}(X,Y)$.
+In TTA, the model is first trained on a labeled source dataset $\{(x_{i},y_{i})\}_{i=1}^{N}$, IID sampled from the source domain.
+In the target dataset $\{x_{i}\}_{i=1}^{M}$, the objective is to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X))\neq Y)$ without accessing the source dataset during adaptation.
 Graph Test-Time Adaptation (GTTA).
 Assume the model consists of a GNN feature encoder $\phi:{\mathcal{X}}\rightarrow{\mathcal{H}}$ and a classifier $g:{\mathcal{H}}\rightarrow{\mathcal{Y}}$.
-The model is trained on source graph ${\mathcal{G}}^{{\mathcal{S}}}=({\mathcal{V}}^{{\mathcal{S}}},{\mathcal{E}}^{{\mathcal{S}}})$ with node labels $\mathbf{y}^{\mathcal{S}}$ and the goal is to enhance model performance on test graph ${\mathcal{G}}^{{\mathcal{T}}}=({\mathcal{V}}^{{\mathcal{T}}},{\mathcal{E}}^{{\mathcal{T}}})$ with distribution shifts that will be defined in Sec.
+The model is trained on source graph ${\mathcal{G}}^{{\mathcal{S}}}=({\mathcal{V}}^{{\mathcal{S}}},{\mathcal{E}}^{{\mathcal{S}}})$ with node labels $\mathbf{y}^{\mathcal{S}}$ and the goal is to enhance model performance on a test graph ${\mathcal{G}}^{{\mathcal{T}}}=({\mathcal{V}}^{{\mathcal{T}}},{\mathcal{E}}^{{\mathcal{T}}})$ with distribution shifts that will be defined in Sec.
 3.1.
-For node classification tasks, we aim to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})$ without accessing ${\mathcal{G}}^{{\mathcal{S}}}$.
+For node classification tasks, we aim to minimize the test error $\varepsilon^{\mathcal{T}}(g\circ\phi)=\mathbb{P}^{\mathcal{T}}(g(\phi(X_{u},{\mathbf{A}}))\neq Y_{u})$ without accessing ${\mathcal{G}}^{{\mathcal{S}}}$ during adaptation.
 2.2 Related Works
-Test-time Adaptation Test-time training (Sun et al., 2020; Liu et al., 2021; Bartler et al., 2022) adapts the source model to the target domain but requires to first add a customized self-supervised losses in model pretraining.
-In contrast, our setup falls into the category of fully test-time adaptation (Wang et al., 2021a; Liang et al., 2024), where we do not alter the model training pipeline.
-Tent (Wang et al., 2021a) adapts the batch normalization parameters by minimizing the entropy, which motivated following-up studies on normalization layer adaptation  (Gong et al., 2022; Zhao et al., 2023; Lim et al., 2023; Niu et al., 2023).
-Some TTA works directly modify the classifier’s prediction, such as LAME (Boudiaf et al., 2022) that directly refines the model’s soft prediction by applying a regularization in the feature space, T3A (Iwasawa & Matsuo, 2021) that classifies test data based on the distance to the pseudo-prototypes derived from pseudo-labels, TAST (Jang et al., 2022) and PROGRAM (Sun et al., 2024) that extend T3A through constructing more reliable prototype graphs.
-However, the above methods are designed for image-based applications and cannot handle the shifts in neighborhood information of graph data.
-Graph Test-time Adaptation Studies on GTTA are in two categories - node and graph classification.
+Test-time Adaptation.
+Test-time training (TTT) approaches necessitate modifications to both the training pipeline and network architecture to accommodate customized self-supervised tasks, such as rotation prediction [33], contrastive learning [34, 35], and clustering [36].
+These requirements, however, can limit their broader applicability.
+In contrast, our proposed setup aligns with the test-time adaptation (TTA) category, which does not alter the original training pipeline [37].
+One prominent TTA method, Tent [28], adapts batch normalization parameters by minimizing entropy.
+This approach has inspired subsequent research focusing on normalization layer adaptation [38, 39, 40, 29].
+Other TTA strategies directly modify the classifier’s predictions.
+For instance, LAME [27] refines the model’s soft predictions by applying regularization in the feature space.
+T3A [26] classifies test data based on its distance to pseudo-prototypes derived from pseudo-labels.
+Building upon T3A, TAST [41] and PROGRAM [42] aim to construct more reliable prototype graphs.
+However, the effectiveness of these methods can be compromised when significant domain shifts occur [43, 44].
+To address larger domain shifts, some methods leverage stored, lightweight source statistics, such as label distributions, to perform test-time alignment [43, 45, 46, 47, 48, 49].
+ActMAD [43], for example, aligns target feature statistics with stored source statistics across multiple layers.
+More recently, FOA [45] focuses on aligning the statistics of classification tokens in vision transformers.
+Note that all the aforementioned methods are primarily designed for image-based applications and may not adequately address shifts in the neighborhood information characteristic of graph data.
+Graph Test-time Adaptation.
+Studies on GTTA are in two categories - node and graph classification.
 Graph classification problems can treat each graph as an i.i.d.
-input, allowing more direct extension of image-based TTA techniques to graphs (Chen et al., 2022; Wang et al., 2022).
-Our work focuses on node classification.
-GTrans (Jin et al., 2023) proposes to augment the target graph at the test time by optimizing a contrastive loss by generating positive views from DropEdge (Rong et al., 2020) and negative samples from the features of the shuffling nodes (Veličković et al., 2019).
-GraphPatcher (Ju et al., 2024) learns to generate virtual neighbors to improve low-degree nodes classification.
-SOGA (Mao et al., 2024) designs a self-supervised loss that relies on mutual information maximization and homophily assumption.
+input, allowing more direct extension of image-based TTA techniques to graphs [50, 51].
+Our work focuses on node classification [30, 31, 32, 52].
+GTrans [30] proposes to augment the target graph at test time by optimizing a contrastive loss by generating positive views from DropEdge [53] and negative samples from the features of the shuffled nodes [54].
+HomoTTT [31] combines self-supervised learning with homophily-based DropEdge and prediction selection.
+SOGA [32] designs a self-supervised loss that relies on mutual information maximization and the homophily assumption.
 These works are mostly built upon heuristics and may not address structure shifts in principle.
 3 Test Error Analysis
-In this section, we characterize the generalization error between source and target graphs and explicitly attribute it to three different kinds of shifts: label shift, feature shift, as well as neighborhood shift.
-Motivated by our theoretical analysis on the generalization error, we then propose TSA algorithm to minimize the across-graph generalization error.
-(a) Reliability of $\gamma$ (b) Source (c) Nbr. Align (d) SNR $\alpha$
+In this section, we characterize the generalization error between the source and target graphs and explicitly attribute it to different types of shifts in graph data, which further motivate our TSA algorithm.
+3.1 Distribution Shifts on Graphs
+Distribution shifts on graphs have been formally studied in previous GDA works [5, 6, 7, 55, 56, 57].
+Following their definitions, we categorize shifts in graphs into two types: feature shift and structure shift.
+For simplicity, our analysis is based on a data generation process: ${\mathbf{X}}\leftarrow{\mathbf{Y}}\rightarrow{\mathbf{A}}$, where the graph structure and node features are both conditioned on node labels.
+Definition 3.1.
+(Feature Shift).
+Assume node features $x_{u},u\in{\mathcal{V}}$ are i.i.d sampled given labels $y_{u}$, then we have $\mathbb{P}({\mathbf{X}}|{\mathbf{Y}})=\prod_{u\in{\mathcal{V}}}\mathbb{P}(X_{u}|Y_{u})$.
+We then define feature shift as $\mathbb{P}^{\mathcal{S}}(X_{u}|Y_{u})\neq\mathbb{P}^{\mathcal{T}}(X_{u}|Y_{u})$.
+Definition 3.2.
+(Structure Shift).
+As the graph structure involves the connecting pattern between nodes including their labels, we consider the joint distribution of the adjacency matrix and labels $\mathbb{P}({\mathbf{A}},{\mathbf{Y}})$, where structure shift, denoted by $\mathbb{P}^{\mathcal{S}}({\mathbf{A}},{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}},{\mathbf{Y}})$, can be decomposed into conditional structure shift (CSS) $\mathbb{P}^{\mathcal{S}}({\mathbf{A}}|{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}}|{\mathbf{Y}})$ and label shift (LS) $\mathbb{P}^{\mathcal{S}}({\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{Y}})$.
+Here, CSS directly reflects the shift in neighborhood information (neighborhood shift) in Fig.
+1.
+3.2 Theoretical Analysis
+Let $\varepsilon^{\mathcal{S}}(g\circ\phi)$ denote the error of a pretrained GNN on the source domain and $\varepsilon^{\mathcal{T}}(g\circ\phi)$ the error of the model when applied to a test graph in the target domain.
+Inspired by Tachet des Combes et al.
+[58], we provide an upper bound on the error gap between the source and target domains, showing how a pretrained GNN (e.g., ERM) can be influenced.
+We denote the balanced error rate of a pretrained node predictor $\hat{Y}_{u}$ on the source domain as $\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\coloneq\max_{j\in\mathcal{Y}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}\neq Y_{u}|Y_{u}=j).$
+Theorem 3.3 (Error Decomposition Theorem (Informal)).
+Suppose ${\mathcal{G}}^{{\mathcal{S}}}$ and ${\mathcal{G}}^{{\mathcal{T}}}$, we can decouple both graphs into independent ego-networks (center nodes and 1-hop neighbors).
+For any classifier $g$ with a GNN encoder $\phi$ in node classification tasks, we have the following upper bound on the error gap between source and target under feature shift and structure shift:
+$\lvert\varepsilon^{\mathcal{S}}(g\circ\phi)-\varepsilon^{\mathcal{T}}(g\circ\phi)\rvert$
+$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2\text{TV}(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}$ $+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Conditional Structure Shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature Shift}}$
+where $TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))$ is the total variation distance between the source and target label distributions.
+and $\Delta_{CE}$ is the error gap that exists if and only if feature shift exists.
+Our bound aligns with the findings of Liu et al.
+[59], which highlight the impact of conditional structure shift, label shift, and feature shift on generalization to the target graph.
+We extend this understanding by deriving an explicit error bound.
+4 Test-Time Structural Alignment (TSA)
+(a) Reliability of $\gamma$ (b) Source (c) Target with Stru. Shift (d) Nbr. Align
 Refer to caption Refer to caption Refer to caption Refer to caption
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 3: (a) Comparison of neighborhood alignment with $\bm{\gamma}$ from model prediction and Oracle on the CSBM graphs (Deshpande et al., 2018).
-The top (or bottom) subfigures represents the assignment under neighbor shift (or neighbor shift plus label shift, respectively).
+Figure 3: (a) Comparison of neighborhood alignment with $\bm{\gamma}$ from model prediction and Oracle on the CSBM graphs [60].
+The top (or bottom) subfigures represent the assignment under neighbor shift (or neighbor shift plus label shift, respectively).
 Nodes are grouped by the entropy of their soft pseudo labels and the y axis shows the accuracy after assigning $\bm{\gamma}$.
@@
 However, the assignment based on pseudo labels is far from optimal (blue).
-From (b) to (c), the figure with the t-SNE visualization of node representations indicates a model trained from the source domain (CSBM with a label distribution $[0.1,0.3,0.6]$) to the target domain (CSBM with a label distribution $[0.3,0.3,0.3]$).
-The color of the nodes represents the ground-truth labels.
-The top subfigures of (b) and (c) show the output given by the GNN encoder while the bottom subfigures show the classifier decision boundaries.
-(d) Analysis of SNR adjustment $\alpha$ with respect to different layers and node degrees on MAG.
-Detail discussed in Sec.
-5.3.
-3.1 Distribution Shifts on Graphs
-Distribution shifts on graphs were formally studied in previous GDA works (Wu et al., 2020; Liao et al., 2021; Zhu et al., 2021; Wu et al., 2022; You et al., 2023; Zhu et al., 2023).
-Following their definition, we categorize shifts in graphs into two types: feature shift and structure shift.
-For simplicity, our analysis is based on a data generation process: ${\mathbf{X}}\leftarrow{\mathbf{Y}}\rightarrow{\mathbf{A}}$, where graph structure and node features are both conditioned on node labels.
-Definition 3.1.
-(Feature Shift).
-Assume node features $x_{u},u\in{\mathcal{V}}$ are i.i.d sampled given labels $y_{u}$, then we have $\mathbb{P}({\mathbf{X}}|{\mathbf{Y}})=\prod_{u\in{\mathcal{V}}}\mathbb{P}(X_{u}|Y_{u})$.
-We then define the feature shift as $\mathbb{P}^{\mathcal{S}}(X_{u}|Y_{u})\neq\mathbb{P}^{\mathcal{T}}(X_{u}|Y_{u})$.
-Definition 3.2.
-(Structure Shift).
-As graph structure involves the connecting pattern between labels, we consider the joint distribution of the adjacency matrix and labels $\mathbb{P}({\mathbf{A}},{\mathbf{Y}})$, where Structure shift, denoted by $\mathbb{P}^{\mathcal{S}}({\mathbf{A}},{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}},{\mathbf{Y}})$, can be decomposed into as conditional structure shift (CSS) $\mathbb{P}^{\mathcal{S}}({\mathbf{A}}|{\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{A}}|{\mathbf{Y}})$ and label shift (LS) $\mathbb{P}^{\mathcal{S}}({\mathbf{Y}})\neq\mathbb{P}^{\mathcal{T}}({\mathbf{Y}})$.
-3.2 Theoretical Analysis
-Let $\varepsilon^{\mathcal{S}}(g\circ\phi)$ denote the error of a pretrained GNN on the source domain and $\varepsilon^{\mathcal{T}}(g\circ\phi)$ the error of the model when applied to a test graph in the target domain.
-Inspired by Tachet des Combes et al.
-(2020), we provide an upper bound on the error gap between source and target domains, showing how a pretrained GNN (e.g., ERM) can be influenced.
-We denote the balanced error rate of a pretrained node predictor $\hat{Y}_{u}$ on the source domain as $\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\coloneq\max_{j\in\mathcal{Y}}\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}\neq Y_{u}|Y_{u}=j).$
-Theorem 3.3 (Error Decomposition Theorem).
-Suppose ${\mathcal{G}}^{{\mathcal{S}}}$ and ${\mathcal{G}}^{{\mathcal{T}}}$, we can decouple both graphs into independent ego-networks (center nodes and 1-hop neighbors).
-For any classifier $g$ with a mean pooling GNN encoder $\phi$ in node classification tasks, we have the following upper bound for on the error gap between source and target under feature shift and structure shift:
-$\lvert\varepsilon^{\mathcal{S}}(g\circ\phi)-\varepsilon^{\mathcal{T}}(g\circ\phi)\rvert$
-$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2\text{TV}(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}$
-$+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Neighborhood shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature shift}}$
-where $TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))$ is the total variation distance between the source and target label distributions.
-and $\Delta_{CE}$ is the error gap that exists if and only if feature shift exists.
-Our bound aligns with the findings of Liu et al.
-(2024), which highlight the impact of neighborhood shift, label shift, and feature shift on generalization to the target graph.
-We extend this understanding by deriving an explicit error bound.
-Notably, neighborhood shift is reduced from conditional structure shift given the assumptions in Thm.3.3.
-4 Test-Time Structural Alignment (TSA)
+Figure (b) to (d) present t-SNE visualizations of node representations during GTTA.
+The model is trained on the source domain in (b) (CSBM with label distribution $[0.1,0.3,0.6]$) and evaluated on the target domain in (c) (CSBM with label distribution $[0.3,0.3,0.3]$).
+Figure (d) illustrates the qualitative result of the neighborhood alignment.
+Despite well-aligned neighborhood representations, a mismatched decision boundary (red) leads to class 0 misclassifications.
+Node colors represent the ground-truth labels.
+The top subfigures show the output given by the GNN encoder, while the bottom subfigures show the classifier decision boundaries.
 Motivated by our theoretical analysis, we propose TSA to address GTTA in a principled way.
-To address neighborhood shift, TSA first conducts neighborhood alignment via weighted aggregation and then properly balances self and neighboring node information combination based on the SNRs of these representations.
-Lastly, as a generic approach for neighborhood shift, TSA can be combined with non-graph TTA methods to refine the decision boundary to address the remaining feature and label shift.
+To address structure shift, TSA first conducts neighborhood alignment via weighted aggregation and then properly balances self and neighboring node information combination based on the SNRs of these representations.
+Lastly, as a generic approach for structure shift, TSA can be combined with non-graph TTA methods to refine the decision boundary.
 4.1 Neighborhood Alignment
-Neighborhood shift alters the node label ratios in the aggregated neighborhood information, causing the shift in the GNN-encoded representations.
-In training-time GDA, PairAlign (Liu et al., 2024) leverages such a technique by assigning edge weights to align the distribution of the source neighborhood with the target domain.
-Inspired by this idea, for GTTA, we aim to achieve a similar goal but in a different direction.
+The conditional structure shift (CSS) alters the node label ratios in the aggregated neighborhood information, causing a shift in the GNN-encoded representations.
 Based on Theorem 3.3, we align the target neighborhood distribution with the source domain, leveraging the fact that the pre-trained model is optimized for the source distribution.
-Specifically, the neighborhood distribution determines the ratio of messages passed from a neighboring class $j$ to center class $i$.
+Specifically, the neighborhood distribution determines the ratio of messages passed from a neighboring class $j$ to the center class $i$.
 To adjust for distributional differences, this ratio can be rescaled by assigning weights to edges from class $j$ to class $i$, effectively acting as an up/down-sampling mechanism during message aggregation.
@@
 $[\bm{\gamma}]_{i,j}=\frac{\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})},\>\forall i,j\in{\mathcal{Y}}$ (2)
-To estimate $\bm{\gamma}$, we assume that the source summary statistics $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ are recorded and available at test time; otherwise, alignment would not be feasible.
-Storing $\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ incurs minimal cost, as it is merely an $|{\mathcal{Y}}|\times|{\mathcal{Y}}|$ matrix.
-Beyond this, no additional information from the source domain is required.
-For $\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$, we estimate it based on target pseudo labels.
-Note that PairAlign (Liu et al., 2024) enhances estimation accuracy that relies pseudo labels (Liu et al., 2023) by leveraging a least-squares constrained optimization.
-However, in GTTA, the absence of source graphs and the potential need for real-time adaptation render this approach impractical.
-While it seems direct to apply a similar alignment strategy by using the inverse of the ratio employed in PairAlign, assigning $\gamma$ requires knowledge of node labels, which is missing in the targe graph, making it challenging in GTTA.
+To estimate $\bm{\gamma}$, we estimate $\mathbb{P}^{\mathcal{T}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ based on target pseudo labels.
+$\mathbb{P}^{\mathcal{S}}(Y_{v}=j|Y_{u}=i,v\in\mathcal{N}_{u})$ can be computed directly from the source graph during model training and added as extra model parameters for the test time direct usage.
+During test time, we weight the message with a properly assigned $[\bm{\gamma}]_{i,j}$ computed based on the ratio of the two quantities.
+This idea aligns with the general principle of test-time alignment [43, 45, 46], but is novel in the aspect of dealing with the structure shift for graph data.
+Assigning edge weights to reduce structure shifts was also studied in training-time GDA by PairAlign [59].
+TSA differs from PairAlign in that $\bm{\gamma}$ estimation and model training can be done together in PairAlign, which better coordinates the two sides as it does training-time GDA (the model better fits the weighted message passing, and the weights are adjusted based on the improved models).
+However, this cannot be done at test time.
+In particular, to compensate for such a loss, TSA needs a more reliable assignment strategy specified in the following.
 Reliable Assignment of $\bm{\gamma}$.
-The ratio $\gamma$ should be assigned to edge weights based on the label pairs of the central and neighboring nodes.
-In training-time GDA, this assignment is straightforward as it relies on the ground-truth labels of the source graph.
-However, in GTTA, this information is unavailable.
-A naive approach is to use target pseudo labels, but this often results in significant mismatches.
-TSA addresses this by quantifying the uncertainty of target pseudo labels (Zhang et al., 2019b; Stadler et al., 2021; Hsu et al., 2022).
+The ratio $\bm{\gamma}$ should be assigned to edge weights based on the label pairs of the central and neighboring nodes.
+In training-time GDA, this assignment is straightforward, as it relies on the ground-truth labels of the source graph.
+However, in GTTA, the node labeling information is unavailable.
+A naive approach is to use target pseudo labels, i.e., $[\bm{\gamma}]_{i,j}$ for the message from a node with pseudo label $j$ to a node with pseudo label $i$, but this often results in significant mismatches.
+TSA addresses this by quantifying the uncertainty of target pseudo labels [61, 62, 63, 64].
 In particular, TSA assigns $[\bm{\gamma}]_{i,j}$ only to node pairs $v\rightarrow u$ where both of their soft target pseudo labels $\hat{y}$ have low entropy $H(\hat{y})=-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$.
-Here, $\rho_{1}$ is a hyperparameter and $\ln(|{\mathcal{Y}}|)$ is the maximum entropy in $|{\mathcal{Y}}|$ class prediction.
-In Fig.3 (a), nodes with low-entropy soft predictions are more reliable, resulting in higher accuracy after the assignment of $\bm{\gamma}$.
+Here, $\rho_{1}$ is a hyperparameter and $\ln(|{\mathcal{Y}}|)$ is the maximum entropy for a $|{\mathcal{Y}}|$ class prediction.
+In Fig.
+3 (a), nodes with low-entropy soft predictions are more reliable, resulting in higher accuracy after the assignment of $\bm{\gamma}$.
+[…]
+Refer to caption Refer to caption Refer to caption
+Figure 4: (a) and (b) compare the SNR of neighborhood-aggregated representations to that of self-node representations across node degrees and GNN layers, based on the ratio of the two quantities.
+A higher ratio indicates that the neighborhood-aggregated representations exhibit higher SNR.
+(c) Analysis of the SNR adjustment $\alpha$ across different layers and node degrees.
+Details are discussed in Sec.
+5.2.
+4.2 SNR-Inspired Adjustment
+Building on neighborhood alignment, we further optimize the signal-to-noise ratio (SNR) of node representations to enhance performance.
+Specifically, the SNR is characterized by the ratio of inter-class representation variance to intra-class representation variance.
+$\text{SNR}\coloneqq\frac{\sigma_{inter}^{2}}{\sigma_{intra}^{2}},\quad\sigma_{inter}^{2}=\frac{1}{|{\mathcal{Y}}|}\sum_{i\in{\mathcal{Y}}}n_{i}\left\|\mu_{i}-\mu_{*}\right\|_{2}^{2},\quad\sigma_{intra}^{2}=\frac{1}{|{\mathcal{Y}}|}\sum_{i\in{\mathcal{Y}}}\sum_{h\in\mathcal{H}_{i}}\left\|h-\mu_{i}\right\|_{2}^{2}$ (3)
+where $n_{i}$ is the number of samples in class $i$, $\mu_{i}$ is the centroid representation of class $i$, $\mu_{*}$ is the overall centroid of all nodes, and $\mathcal{H}_{i}$ represents the set of representations in class $i$.
+A higher inter-class variance $\sigma_{inter}^{2}$ indicates informative and well-separated representations, while a higher $\sigma_{intra}^{2}$ indicates a higher noise in the representations.
+Optimizing SNR complements the neighborhood alignment approach.
+Even if neighborhood label distributions are perfectly aligned, variations in neighbor set sizes between the source and target graphs can impact the SNR of aggregated neighboring node representations.
+To compare the SNR of neighborhood-aggregated representations to that of self-node representations, we compute SNR ratio=Nbr.
+Aggr.
+SNRSelf-Node $\text{SNR ratio}=\frac{\text{Nbr.
+Aggr.
+SNR}}{\text{Self-Node SNR}}$ which exhibits different trends as a function of node degree (aka neighbor set size) and layer depth, as shown in Fig.
+4 (a) and (b).
+In both figures, deeper layers and higher degree values generally lead to a higher SNR ratio, indicating aggregated representations are better denoised.
+However, the rate and shape of SNR ratio growth vary across source and target domains.
+Consequently, the combination of these aggregated representations with self-node representations in a typical GNN pipeline (Eq.
+1) should be properly adjusted as well in test-time adaptation.
+To implement the SNR-inspired adjustment, we introduce a parameter to perform the weighted combination of self-node representations and neighborhood-aggregated representations at each layer, adapting to node degrees as follows:
+Definition 4.2.
+Let $\tilde{d_{u}}=\left(\frac{\ln(d_{u}+1)}{\ln(\max_{v\in\mathcal{V}}d_{v}+1)}\right)$ denote log-normalized degree of node $u$ and let $\text{MLP}^{(k)}$ and $b^{(k)}$ to be learnable parameters for adjusting $k$-th layer combination, define the weights for combination $\alpha\in\mathbb{R}^{L}$ as:
+$[\alpha]_{k}=\sigma(\text{MLP}^{(k)}(\tilde{d_{u}}))-0.5+b^{(k)},\>\forall k\in[L]$ (4)
+where $\sigma$ is the sigmoid function. During initialization $[\alpha]_{k}$ is set to $1$. Degree values are taken in the logarithmic domain to handle their often long-tailed distribution.
+Together, $[\alpha]_{k}\cdot[\bm{\gamma}]_{i,j}$ is used to reweight the GNN messages, adjusting the influence from the node with a highly certain pseudo label $j$ to the node with a highly certain pseudo label $i$.
+Remark.
+Neighborhood alignment alone does not address potential shifts in SNR.
+This is because the alignment approach, inspired by Thm.
+3.3, focuses solely on population-level performance, whereas SNR also incorporates the noise (variance) induced by empirical sampling from the population.
+Thus, these two aspects complement each other.
 Algorithm 1 Test-Time Structural Alignment (TSA)
@@
 9:  return $\hat{y}_{\mathrm{final}}$ after another boundary refinement
-4.2 SNR-Inspired Adjustment
-Building on neighborhood alignment, we further optimize the signal-to-noise ratio (SNR) of node representations to enhance performance.
-Specifically, SNR is characterized by the ratio of inter-class representation distances to intra-class representation distances.
-A higher SNR indicates more informative and well-separated representations, which benefit classification.
-Optimizing SNR complements the neighborhood alignment approach.
-Even if neighborhood label distributions are perfectly aligned, variations in neighbor set sizes between the source and target graphs can impact the SNR of aggregated neighboring node representations.
-Consequently, the combination of these aggregated representations with self-node representations in a typical GNN pipeline (Eq.
-1) should be adjusted accordingly across source and target domains, particularly when the source and target graphs exhibit very different degree distributions.
-Furthermore, the SNR of node self-representations may vary across GNN layers, as deeper layers generally reduce variance.
-As a result, node self-representations in deeper layers tend to have higher SNR and should be assigned greater emphasis.
-To implement the SNR-inspired adjustment, we introduce a parameter to perform the weighted combination of self-node representations and neighborhood-aggregated representations at each layer, adapting to node degrees as follows:
-Definition 4.2.
-Let $\tilde{d_{u}}=\left(\frac{\ln(d_{u}+1)}{\ln(\max_{v\in\mathcal{V}}d_{v}+1)}\right)$ denote log-normalized degree of node $u$ and let $\text{MLP}^{(k)}$ and $b^{(k)}$ to be learnable parameters for adjusting $k$-th layer combination, define the weights for combination $\alpha\in\mathbb{R}^{L}$ as:
-$[\alpha]_{k}=\sigma(\text{MLP}^{(k)}(\tilde{d_{u}}))-0.5+b^{(k)},\>\forall k\in[L]$ (3)
-where $\sigma$ is the sigmoid function. During initialization $[\alpha]_{k}$ is set to $1$. Degree values are taken in the logarithmic domain to handle their often long-tailed distribution.
-Combined with $\bm{\gamma}$, $[\alpha]_{k}\cdot[\bm{\gamma}]_{i,j}$ is used to reweight the GNN message for non-self-loop node pairs, adjusting the influence from the node with a highly certain pseudo label $j$ to the node with a highly certain pseudo label $i$.
-Remark.
-Neighborhood alignment alone does not address potential shifts in SNR.
-This is because the alignment approach, inspired by Thm.
-3.3, focuses solely on expectation-based (first-order statistical) performance, whereas SNR also incorporates variance (second-order statistics).
-Thus, these two aspects complement each other.
 4.3 Decision Boundary Refinement
-In GTTA, label shift can result in a mismatch of the decision boundary, even after addressing neighborhood shift and obtaining high-SNR node representations. This is illustrated in Fig. 3 (b) and (c).
-A straightforward approach to refining the decision boundary at test time is to adjust the classifier’s batch normalization parameters (TENT (Wang et al., 2021a)) or directly modify its output (T3A (Iwasawa & Matsuo, 2021) and LAME  (Boudiaf et al., 2022)).
+In GTTA, even when CSS is fully addressed and high-SNR node representations are obtained, the decision boundary may still suffer from a mismatch due to label and feature shifts (Fig. 3 (d)).
+A straightforward approach to refining the decision boundary at test time is to adjust the classifier’s batch normalization parameters (TENT [28]) or directly modify its output (T3A [26] and LAME  [27]).
 We integrate these techniques into our framework for two folds: (1) their refined pseudo-labels provide a more reliable assignment of $\bm{\gamma}$ and can supervise the update of SNR adjustment.
 (2) Reciprocally, better alignment of neighborhood information can further refine the decision boundary.
-4.4 TSA Overview.
+Matcha [65] is a very recent work on GTTA that also integrates non-graph TTA methods with an approach for addressing structure shift.
+However, it considers only degree shift and homophily shift, where homophily shift is merely a special case of CSS in our context.
+Consequently, Matcha does not introduce the parameter $\bm{\gamma}$ and does not fully address the structure shift.
+TSA Overview.
 Note that $\bm{\gamma}$ is obtained from the initial pseudo labels.
 Only the parameters in $\text{MLP}^{(k)}$ and $b^{(k)}$ for $\alpha$ estimation in Eq.
-3 need to be optimized in the test time according to Eq.
-4.
+4 need to be optimized at test time according to Eq.
+5.
 ${\mathcal{L}}_{CE}=\frac{1}{|{\mathcal{V}}^{{\mathcal{T}}}|}\sum_{u\in{\mathcal{V}}^{{\mathcal{T}}}}\text{cross-entropy}(y_{u}^{\prime},\hat{y}_{u})$ (5)
-where $y_{u}^{\prime}$ is the hard pseudo label refined by procedure in Sec.
+where $y_{u}^{\prime}$ is the hard pseudo label refined by the procedure in Sec.
 4.3 and $\hat{y}_{u}$ is the soft prediction from the original model.
@@
 4.3.
-Our proposed algorithm is summarized in Alg.
+Our algorithm is summarized in Alg.
 1.
-Comparison to AdaRC (Bao et al., 2024) AdaRC is a recent work on GTTA that also integrates non-graph TTA methods with an approach for addressing structure shifts.
-However, it considers only degree shift and homophily shift, where homophily shift is merely a special case of neighborhood shift in our context.
-Consequently, AdaRC does not introduce the parameter $\bm{\gamma}$ and, therefore, lacks the ability to properly align neighborhood aggregated representations when the neighboring label distribution shifts.
 5 Experiments
-Table 1: MAG results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+We evaluate TSA on synthetic datasets and 5 real-world datasets. Additional results and discussions (e.g., hyperparameter sensitivity) are provided in Appendix D.
+5.1 Datasets and Baselines
+Synthetic Data.
+We use the CSBM model [60] to generate three-class datasets, focusing on structure shifts while keeping feature distributions unchanged.
+Specifically, we evaluate performance under three conditions: (1) CSS, (2) CSS plus SNR (induced by degree changes) shift, and (3) full structure shift (CSS + label shift) plus SNR shift.
+This experimental setup is motivated by Thm.
+3.3 and the observations in Sec.4.2.
+For each shift scenario, we examine two levels of severity, with the left column in Table 1 corresponding to the smaller shift.
+MAG [59] is a citation network extracted by OGB-MAG [66].
+Distribution shifts arise from partitioning papers into distinct graphs based on their countries of publication.
+The task is to classify the publication venue of each paper.
+Our model is pretrained on graphs from the US and China and subsequently adapted to graphs from other countries.
 […]
+Arxiv [66] is a citation network between all Computer Science (CS) arXiv papers.
+Distribution shifts originate from different time periods.
+Our model is pretrained on the earlier time span 1950 to 2007/ 2009/ 2011 and tested on later periods 2014-2016 and 2016-2018.
+DBLP and ACM [68, 5] are two citation networks. The model is trained on one network and adapted to the other to predict the research topic of a paper (node).
+Baselines.
+We compare TSA with six baselines.
+For non-graph TTA methods, we include TENT [28], LAME [27], T3A [26], and ActMAD [43].
+Note that TENT is applied only to the classifier, as GNNs typically do not include batch normalization layers.
+GTrans [30], HomoTTT [31], SOGA [32], and Matcha [65] are direct comparisons in GTTA.
+Matcha is limited to GPRGNN due to its design.
+We present the results for GraphSAGE [22] in the main paper, while the results for GPRGNN [69] (including Matcha) and GCN [20] are provided in the Appendix.
+5.2 Result Analysis
 Table 1: Synthetic CSBM results (accuracy).
@@
 GTrans 86.67$\pm$3.59 72.37$\pm$4.06 79.55$\pm$1.23 68.69$\pm$3.27 59.24$\pm$2.12 47.42$\pm$3.73 79.29$\pm$2.71 66.77$\pm$2.52
+HomoTTT 85.08$\pm$2.87 65.79$\pm$9.07 78.84$\pm$3.42 65.31$\pm$5.69 52.34$\pm$4.69 40.68$\pm$5.24 70.11$\pm$5.08 62.19$\pm$2.51
 SOGA 86.09$\pm$3.89 70.39$\pm$7.96 79.75$\pm$3.20 69.00$\pm$5.28 56.60$\pm$3.88 44.09$\pm$5.32 73.52$\pm$5.30 63.76$\pm$3.45
+ActMAD 84.95$\pm$4.04 65.17$\pm$10.19 78.66$\pm$3.70 64.95$\pm$5.99 54.78$\pm$3.94 42.22$\pm$4.83 72.41$\pm$5.67 63.39$\pm$3.33
 TENT 87.48$\pm$2.86 77.14$\pm$4.64 81.04$\pm$2.72 72.51$\pm$3.39 76.48$\pm$6.21 59.12$\pm$5.06 77.21$\pm$5.53 62.36$\pm$6.83
@@
 TSA-T3A 89.96$\pm$1.33 81.08$\pm$2.73 84.23$\pm$1.24 76.89$\pm$2.02 65.59$\pm$2.57 52.34$\pm$7.19 82.55$\pm$5.06 64.88$\pm$3.30
+Table 2: MAG results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+[…]
 Table 3: Pileup results (f1-scores). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
@@
 GTrans 57.37$\pm$1.49 63.66$\pm$2.43 48.13$\pm$2.11 65.74$\pm$1.95 28.41$\pm$4.01 57.65$\pm$1.97 69.17$\pm$0.82 67.37$\pm$0.56
+HomoTTT 58.29$\pm$1.34 62.50$\pm$3.35 49.07$\pm$1.32 65.18$\pm$1.76 28.63$\pm$2.34 56.11$\pm$3.83 70.31$\pm$0.58 68.85$\pm$0.29
 SOGA 60.13$\pm$0.75 69.71$\pm$0.92 51.68$\pm$0.83 67.47$\pm$1.42 37.16$\pm$1.15 56.59$\pm$3.86 70.83$\pm$0.66 68.84$\pm$0.97
+ActMAD 58.54$\pm$0.81 64.59$\pm$2.38 49.39$\pm$1.35 67.12$\pm$1.74 20.88$\pm$2.66 50.97$\pm$5.14 69.20$\pm$0.79 68.04$\pm$0.46
 TENT 58.25$\pm$1.85 58.73$\pm$4.80 48.27$\pm$2.58 58.41$\pm$7.89 30.59$\pm$3.07 0.04$\pm$0.08 68.72$\pm$0.48 68.00$\pm$0.60
@@
 TSA-T3A 61.03$\pm$1.19 70.42$\pm$1.66 52.39$\pm$0.86 70.92$\pm$0.64 37.33$\pm$2.59 57.73$\pm$1.98 69.73$\pm$1.09 68.75$\pm$0.66
-Table 4: Arxiv and DBLP/ACM no balanced training (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+Table 4: Arxiv and DBLP/ACM results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
 1950-2007 1950-2009 1950-2011 DBLP & ACM
@@
 GTrans 40.92$\pm$0.32 40.25$\pm$1.69 45.31$\pm$1.99 43.83$\pm$3.15 53.68$\pm$0.95 52.57$\pm$1.23 34.47$\pm$2.99 49.68$\pm$5.44
+HomoTTT 36.33$\pm$0.95 32.83$\pm$1.15 41.22$\pm$1.15 38.86$\pm$1.42 48.18$\pm$1.38 44.61$\pm$2.90 31.25$\pm$3.96 52.90$\pm$6.25
 SOGA 34.11$\pm$2.91 28.94$\pm$4.68 41.59$\pm$2.03 39.61$\pm$3.22 50.12$\pm$1.38 43.03$\pm$3.70 35.74$\pm$5.20 58.59$\pm$8.35
+ActMAD 41.02$\pm$0.53 40.43$\pm$1.44 44.88$\pm$1.94 42.39$\pm$3.45 53.43$\pm$1.13 51.78$\pm$1.78 29.47$\pm$4.46 52.75$\pm$6.89
 TENT 40.77$\pm$0.32 39.58$\pm$1.30 45.74$\pm$1.26 43.98$\pm$1.90 54.45$\pm$1.03 53.19$\pm$1.31 36.59$\pm$4.26 53.38$\pm$6.60
@@
 TSA-T3A 40.03$\pm$1.03 39.70$\pm$1.49 44.09$\pm$1.53 42.17$\pm$2.91 50.96$\pm$1.68 49.21$\pm$2.67 35.30$\pm$5.57 52.55$\pm$6.92
-We evaluate TSA on synthetic datasets and 5 real-world datasets. More discussions and results in experiments can be found in Appendix D.
-5.1 Datasets and Baselines
-Synthetic Data.
-We use the CSBM model (Deshpande et al., 2018) to generate three-class datasets, focusing on structure shifts while keeping feature distributions unchanged.
-Specifically, we evaluate performance under three conditions: (1) neighborhood shift, (2) neighborhood shift plus SNR (induced by degree changes) shift, and (3) neighborhood shift combined with both SNR shift and label shift.
-This experimental setup is motivated by Thm.
-3.3 and the observations in Sec.4.2.
-For each shift scenario, we examine two levels of severity, with the left column in later Table 2 corresponding to the smaller shift.
-MAG (Liu et al., 2024) is a citation network extracted by OGB-MAG (Hu et al., 2020).
-Distribution shifts arise from partitioning papers into distinct graphs based on their countries of publication.
-The task is to classify the publication venue of each paper.
-Our model is pretrained on graphs from the US and China and subsequently adapted to graphs from other countries.
-[…]
-Arxiv (Hu et al., 2020) is a citation network between all Computer Science (CS) arXiv papers.
-Distribution shifts originate from different time.
-Our model is pretrained on the earlier time span 1950 to 2007/ 2009/ 2011 and test on later 2014 to 2016 and 2016 to 2018.
-DBLP and ACM (Tang et al., 2008; Wu et al., 2020) are two citation networks. The model is trained on one network and adapted to the other to predict the research topic of a paper (node).
-Baselines We compare TSA with six baselines.
-For non-graph TTA methods, we include TENT (Wang et al., 2021a), LAME (Boudiaf et al., 2022), and T3A (Iwasawa & Matsuo, 2021).
-Note that TENT is only applied to the classifier, as GNNs typically do not include batch normalization layers.
-GTrans (Jin et al., 2023), SOGA (Mao et al., 2024), and AdaRC (Bao et al., 2024) are direct comparisons in GTTA.
-AdaRC is limited to GPRGNN due to its design.
-We present the results for GraphSAGE (Hamilton et al., 2017) in the main paper, while the results for GPRGNN (Chien et al., 2020) (including AdaRC) are provided in the Appendix.
-5.2 Result Analysis
-From the MAG dataset results in Table 1, TSA combined with boundary refinement techniques consistently achieves top performance among all baselines.
-Compared to non-graph TTA baselines, such as TENT vs.
-TSA-TENT, TSA also provides considerable improvements over the original adaptation results.
-Notably, the benefits of TSA after neighborhood alignment are more pronounced in scenarios with larger neighborhood shifts, such as adaptation from the US to FR, as reflected in the dataset shift metrics (Table LABEL:table:magstats).
-Among all baselines, T3A consistently outperforms as a representative non-gradient-based method that refines the decision boundary, while TENT struggles in most scenarios.
-This is primarily due to the relatively imbalanced class distribution in the MAG datasets, where evaluation is conducted on the top 19 classes.
-Since TENT learns feature-wise transformations through entropy optimization on the target graph, it tends to be biased toward the dominant class.
-Other GTTA methods, such as GTrans, achieve performance comparable to ERM, whereas SOGA sometimes underperforms due to its strong reliance on the graph homophily assumption, which is often weak in sparse graphs with many classes, such as MAG and Arxiv.
-The results for the synthetic datasets are presented in Table 2.
-The first six columns correspond to experiments conducted under imbalanced source training.
-Compared to real datasets, we observe more significant advantages from TSA.
-TENT outperforms LAME and T3A, as it directly maximizes the most probable class through entropy minimization.
-In contrast, LAME incorporates regularization, while T3A relies on distance to prototypes, making it less sensitive to majority-class dominance (Zhang et al., 2022).
-When combined with TSA, we observe similar results across all three variants, indicating that the primary adaptation ability stems from TSA rather than non-graph TTA methods.
-While GTTA baselines perform better than ERM, they still fall short of our TSA-based methods.
-TSA also demonstrates strong performance under additional shifts in SNR caused by degree shifts.
-Additionally, we consider training a pretrained model on a balanced source domain under the same edge connection probability as the previous two columns.
-The last two columns demonstrate better performance than the previous two under balanced training, even though they encounter the same shift in label connection.
-This showcases that dataset imbalance is a severe issue in source training, which aligns with the worst-case error described in Theorem 3.3.
-The pileup results in Table 3 demonstrate the effectiveness of TSA-based methods under both neighborhood shift and label shift, with each module contributing in different ways.
-The neighborhood alignment module provides significant improvements when generalizing from a low PU level to a high PU level, as this scenario is primarily dominated by neighborhood shift.
-Conversely, when adapting from a high PU level to a low PU level, label shift becomes the dominant factor.
-In this case, incorporating non-graph TTA techniques is crucial, while neighborhood alignment offers an additional but smaller benefit.
-Moreover, TENT can be vulnerable when generalizing from an imbalanced source graph to a balanced target graph.
-This is because source training on high PU graphs suffers from class imbalance, making entropy minimization non-robust due to a mismatched decision boundary when adapting to low PU graphs.
-The last two columns, which represent the same PU level but different physical signals, show minimal benefits from TSA.
-This is expected, as these scenarios primarily exhibit feature shifts rather than structure or label shifts.
-The results from Arxiv and DBLP/ACM in Table 4 confirm that TSA can be effectively integrated with non-graph TTA methods to mitigate feature shift while also facilitating reasonable adjustments for structure shifts.
-These citation networks exhibit relatively mild structure shifts, leading to closely comparable performance across all baselines.
-TENT and TSA-TENT stand out by maximizing the majority classes through entropy minimization.
-Nonetheless, we still observe meaningful improvements with the addition of TSA when compared to the corresponding baselines, demonstrating its effectiveness in enhancing adaptation.
-5.3 Analysis of $\alpha$ for SNR-induced refinement
-As defined in Eq.3, $\alpha$ is learned, depending on node degrees and GNN layer depth.
+Comparison with Baselines.
+TSA achieves strong performance across all datasets and GNN backbones.
+In all cases, the best results are obtained by GTTA methods (underlined across Tables 14), highlighting the importance of addressing structure shift in graph data.
+In the synthetic dataset (Table 1), TSA outperforms GTTA baselines under plain CSS, demonstrating neighborhood alignment as an effective approach to handling CSS.
+The same table also showcases that dataset imbalance severely affects source training, which aligns with the worst-case error described in Theorem 3.3.
+In real-world datasets, TSA achieves an average improvement of 10% over existing GTTA baselines.
+GTrans relies on self-supervised learning to augment graphs.
+Although it shows improvements on CSBM, its performance is limited on real-world datasets (Tables 24).
+SOGA and HomoTTT depend on the homophily assumption, which does not always hold in real-world graphs [70], leading to unstable performance.
+In contrast, TSA makes no such assumptions and consistently delivers stable improvements by structure shifts.
+Different TSA-version.
+Overall, TSA-variants consistently yield performance gains of up to 21.25% over the corresponding non-graph TTA baselines (bold across Table 1- 4).
+In particular, the performance of the TSA variants differs in how the boundary refinement addresses the label shift.
+TSA-T3A is less sensitive to the dominance of the majority class, as it refines the boundary based on distances to prototypes, offering increased robustness [71].
+Therefore, TSA-T3A works well on imbalanced datasets such as MAG (Table 2) and Pileup (Table 3).
+We observe TSA-TENT and TSA-LAME perform well on Arxiv and ACM/DBLP (Table 4), as they directly maximize the most probable class, with TSA-LAME introducing a regularization overhead.
+Analysis of $\alpha$.
+The learned $\alpha$ depends on node degrees and GNN layer depth.
 The trend of $\alpha$ in Fig.
-3 (d) aligns with our expectations: earlier GNN layers require less attention to self-node representations (small $\alpha$), while deeper layers increasingly rely on them.
-Additionally, the plot reveals a correlation with node degrees, where nodes with higher degrees (and thus higher SNR) place greater weight on neighborhood-aggregated messages.
+4 (c) aligns with our expectations: earlier GNN layers require less attention to neighborhood-aggregated representations (small $\alpha$), while deeper layers increasingly rely on them.
+Additionally, the plot reveals a correlation with node degrees, where nodes with higher degrees (and thus higher SNR ratios) place greater weight on neighborhood-aggregated messages.
 However, the overall effect of different GNN layers appears to be greater than that of varying node degrees.
-5.4 Ablation Studies
-In Table 6 (Appendix), we examine the effects of neighborhood alignment and SNR-based representation adjustments.
+Lastly, the learned $\alpha$ exhibits a trend similar to the source SNR ratio shown in Fig.
+4 (a), assigning lower weights in layer 1 and nearly identical weighting patterns in layer 2 and 3.
+Ablation Studies.
+In Table 5 (Appendix), we examine the gains of neighborhood alignment and SNR-based representation adjustments on both synthetic and real-world datasets.
 The results consistently show that incorporating both modules yields the best performance.
-In some cases, removing SNR adjustments leads to a greater performance drop compared to removing neighborhood alignment.
-However, this does not necessarily indicate that neighborhood alignment is less important.
-Instead, we think this effect arises because the additional backpropagation training for SNR adjustments amplifies their contributions.
-6 Discussion and Limitations
+Removing neighborhood alignment leads to a performance drop of up to 19.9% on the synthetic dataset and 3.16% on the real-world dataset, while removing SNR adjustments results in a drop of up to 2.43% and 2.97%, respectively.
+Removing neighborhood alignment causes a larger drop in CSBM due to the more significant shift used in the data modeling.
+Computational Efficiency.
+In Table 6 (Appendix), we quantify the computation time of TSA-T3A on MAG US, containing the largest number of nodes and edges.
+The time complexity of neighborhood alignment is $O(1)$ as it is a one-step edge weight assignment.
+SNR adjustment contributes relatively more to the overhead due to backpropagation.
+Notably, compared to GTTA baselines, TSA-T3A is approximately 2, 4 and 8 times more efficient than HomoTTT, GTrans and SOGA.
+6 Conclusion
 In this work, we theoretically analyze the impact of various graph shifts on the generalization gap in GTTA and highlight their empirical effects on test-time performance degradation.
-Based on these insights, we propose the TSA algorithm, which addresses neighborhood shift through extended and improved neighborhood alignment strategies tailored to the unique challenges of GTTA, and handles label shift via classifier adjustments.
-Despite its effectiveness, we observe that TSA ’s benefits can be marginal when feature shift dominates, as evidenced in the last two columns of Pileup, Arxiv, and DBLP/ACM.
-This is primarily because the neighborhood alignment strategy is not robust in such scenarios.
-Moreover, our current estimation of $\bm{\gamma}$ heavily depends on the accuracy of refined soft pseudo-labels.
-While confusion-matrix-based methods (Lipton et al., 2018; Azizzadenesheli et al., 2019; Alexandari et al., 2020) can provide more robust estimations, they often require additional adversarial training, which is not compatible with GTTA constraints.
-We encourage future work to explore more robust approaches for estimating $\bm{\gamma}$ to further enhance adaptation under severe feature shifts.
-7 Broader Impact
-This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
+Based on these insights, we propose TSA, a method that addresses structure shift through neighborhood alignment, SNR-inspired adjustment, and boundary refinement.
+Extensive experiments validate the effectiveness and efficiency of TSA.
 7 Acknowledgment
@@
 $=\sum_{i\neq j}\gamma^{\mathcal{U}}_{u_{j}}\sum_{k\in\mathcal{Y}}\pi^{\mathcal{U}}_{v_{k}|u_{j}}\omega^{\mathcal{U}}_{u_{i}|u_{j},v_{k}}$
-The function of mean pooling naturally marginalizes the effect of degree magnitude, meaning that only the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself influence the prediction.
+Most GNNs marginalize the effect of degree magnitude, meaning that only the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself influence the prediction.
 We use Lemma A.1 to bound the terms above.
@@
 (a) comes from Lemma A.1.
-(b) comes from setting $\alpha_{j}=1$ and $\beta_{j}=0,\forall j\in\mathcal{Y}$ (c) The inequality comes from Holder’s inequality, where we take the maximum-norm for $k$ and take the $L_{1}$ norm to marginalize $Y_{v}$: $\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)=\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ (d) The last term in the previous line corresponds to the conditional error gap in (Tachet des Combes et al., 2020) but in a GNN mean pooling fashion.
-The term $\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ can be interpreted as the probability of error prediction using a mean pooling GNN encoder.
-By definition, $Y_{v}$ represents a random variable of the neighborhood label ratio and $Y_{u}$ denotes a random variable of self class label, their realization corresponds to the mean pooled and center value.
-As a reult, $\Delta_{CE}$ only exists if and only if feature shift exists.
+(b) comes from setting $\alpha_{j}=1$ and $\beta_{j}=0,\forall j\in\mathcal{Y}$ (c) The inequality comes from Holder’s inequality, where we take the maximum-norm for $k$ and take the $L_{1}$ norm to marginalize $Y_{v}$: $\mathbb{P}^{\mathcal{S}}(\hat{Y}_{u}=i|Y_{u}=j)=\sum_{k\in\mathcal{Y}}\mathbb{P}^{\mathcal{U}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ (d) The last term in the previous line corresponds to the conditional error gap in [58] but in a standard GNN’s message passing fashion.
+The term $\mathbb{P}^{\mathcal{U}}(\hat{Y}_{u}=i|Y_{u}=j,Y_{v}=k,v\in\mathcal{N}_{u})$ can be interpreted as the probability of error prediction using a GNN encoder.
+By definition, $Y_{v}$ represents a random variable of the neighborhood label ratio and $Y_{u}$ denotes a random variable of self class label, their realization corresponds to the expectation of the neighboring nodes’ label distribution and the label distribution of the central node itself.
+As a result, $\Delta_{CE}$ only exists if and only if feature shift exists.
 Putting everything together, we have
 $\lvert\varepsilon^{\mathcal{S}}(g\circ\phi)-\varepsilon^{\mathcal{T}}(g\circ\phi)\rvert$
-$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Neighborhood shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature shift}}\vspace{-2mm}$
-Appendix B Additional Results
-B.1 Results of GPRGNN
-Table 5 presents the results of the GPRGNN backbone on the MAG dataset.
-Compared to GraphSAGE, we observe that GPRGNN is less robust under structure shifts due to its linear combination of k-hop representations.
-We also notice a similar tendency as with GraphSAGE, where T3A outperforms all other baselines.
-TSA slightly improves upon T3A by refining the decision boundary.
-However, we frequently observe a performance decrease when combining AdaRC with T3A.
-This drop is attributed to the fact that their proposed PIC loss relies on minimizing intra-variance, which depends on the quality of node hidden representations on the target domain.
-Under imbalanced training and label shift, node hidden representations can easily degrade, causing AdaRC performance to become unstable.
-Table 5: MAG GPRGNN results (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+$\leq\textnormal{BER}^{\mathcal{S}}(\hat{Y}_{u}\;\|\;Y_{u})\cdot\{\underbrace{(2TV(\mathbb{P}^{\mathcal{S}}(Y_{u}),\mathbb{P}^{\mathcal{T}}(Y_{u}))}_{\text{Label Shift}}+\underbrace{\mathbb{E}_{Y_{u}}^{\mathcal{T}}[\max_{k\in{\mathcal{Y}}}\lvert 1-\frac{\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}{\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u},v\in\mathcal{N}_{u})}\rvert]}_{\textbf{Conditional Structure Shift}}\}+\underbrace{\Delta_{CE}}_{\text{Feature Shift}}\vspace{-2mm}$
+Appendix B Limitations
+While TSA can address feature shift to some extent, we observe that TSA ’s benefits can be marginal when feature shift dominates significantly over structure shift, as evidenced in the last two columns of Pileup.
+This is primarily because the neighborhood alignment strategy is not robust in such scenarios.
+Moreover, our current estimation of $\bm{\gamma}$ heavily depends on the accuracy of refined soft pseudo-labels.
+While confusion-matrix-based methods [72, 73, 74] can provide more robust estimations, they often require additional adversarial training, which is not compatible with GTTA constraints.
+We encourage future work to explore more robust approaches to estimating $\bm{\gamma}$ to further enhance adaptation under severe feature shifts.
+Appendix C Broader Impact
+Our work studies structure shift and proposes TSA for graph test-time adaptation (GTTA).
+The goal of GTTA is to enhance model performance under distribution shifts with computation and privacy constraints.
+We do not foresee any direct negative societal impacts.
+Appendix D Additional Results
+D.1 TSA Ablation Studies
+In the ablation studies, we examine the effects of neighborhood alignment and SNR-based representation adjustments.
+We select synthetic datasets with structure shift, as well as real-world datasets that exhibit the most conditional structure shift (CSS), according to Table LABEL:table:magstats.
+A detailed discussion of the ablation results is provided in Section 5.2.
+Table 5: TSA Ablation studies on CSBM and MAG.
+[…]
+TSA-TENT Base 88.68$\pm$4.99 66.25$\pm$7.75 32.53$\pm$0.91 28.62$\pm$0.63
+w/o Nbr. Align 78.45$\pm$5.51 61.12$\pm$4.86 32.26$\pm$0.95 28.41$\pm$0.59
+w/o SNR 86.93$\pm$5.05 63.82$\pm$8.06 32.52$\pm$0.90 28.57$\pm$0.59
+TSA-LAME Base 65.09$\pm$2.34 52.89$\pm$6.11 32.86$\pm$2.23 27.22$\pm$1.48
+w/o Nbr. Align 48.45$\pm$5.30 37.31$\pm$3.95 32.70$\pm$2.29 26.49$\pm$1.76
+w/o SNR 63.65$\pm$2.65 51.18$\pm$5.84 29.89$\pm$1.94 24.41$\pm$1.43
+TSA-T3A Base 65.59$\pm$2.57 52.34$\pm$7.19 46.61$\pm$0.88 43.45$\pm$0.81
+w/o Nbr. Align 45.69$\pm$5.44 36.75$\pm$3.09 46.44$\pm$0.80 42.09$\pm$1.26
+w/o SNR 64.17$\pm$2.11 51.04$\pm$6.79 46.51$\pm$0.94 41.35$\pm$1.09
+D.2 Visualization of Distribution Shift
+Figure 5 (a)-(e) investigates the impact of different types of distribution shifts on GNNs.
+We use a one-layer GraphSAGE followed by an MLP classifier to visualize node representations on synthetic datasets generated by the CSBM model.
+  • Feature Shift: Figure 5 (b) shows that the representations deviate from the source domain, but the clusters remain separable.
+  • Conditional Structure Shift (CSS): Figure 5 (c) shows that the embeddings largely overlap, making the representation space less discriminative.
+  • Label Shift: Figure 5 (d) illustrates a biased decision boundary, which leads to misclassification of the minority class.
+  • SNR Shift: Figure 5 (e) shows that the node representations remain roughly aligned with the source, but are more dispersed, as the SNR shift introduces noise into the representations.
+(a) Source (b) Feature Shift (c) CSS (d) Label Shift (e) SNR Shift
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
+Figure 5: The t-SNE visualization of a one-layer GraphSAGE under different distribution shifts: (top) the output representations given by the GNN encoder and (bottom) the node representations with the classifier decision boundaries.
+(a) indicates the source domain where the model is pretrained.
+(b), (c), and (d) stand for feature shift, conditional structure shift (CSS), and label shift.
+(e) represent the impact of SNR shift induced by degree changes.
+The color of the nodes represents the ground-truth labels.
+The legends in the bottom plots show the accuracy and recall scores for each class.
+The red contours in the bottom (d) highlight the decision boundary of the minority class.
+Table 6: Computation time on MAG (CN $\rightarrow$ US). TTA methods are recorded in adaptation time per epoch. The Wall-Clock Time (seconds) are measured on a single NVIDIA RTX A6000 GPU.
+Method Stage Computation Time (sec) Additional Time (%) Inference 0.518$\pm$0.038 TSA-T3A Adaptation 1.138$\pm$0.072 219.69% - Neighborhood Alignment 0.037$\pm$0.011 7.14% - SNR Adjustment 0.729$\pm$0.023 140.73% - Boundary Refinement (T3A) 0.368$\pm$0.046 71.04% GTrans Adaptation 4.309$\pm$0.171 831.85% HomoTTT Adaptation 2.014$\pm$0.815 388.80% SOGA Adaptation 9.264$\pm$0.218 1788.42%
+D.3 Computational Efficiency
+We quantify the computation time of TSA when adapting GraphSAGE on the MAG US dataset and compare it with other GTTA methods:
+  • Inference time refers to the time required for the model to generate predictions on the target graph ${\mathcal{G}}^{{\mathcal{T}}}$. This is also equivalent to the computation time for ERM.
+  • Adaptation time refers to the time required per epoch after inference. It reflects the additional overhead introduced by the adaptation methods.
+For a fair comparison, we report the computation time per epoch of adaptation.
+Note that GTrans alternates between optimizing adjacency structure and node features; in this case, we report the combined runtime of one graph feature update and one edge update as a single adaptation epoch.
+Table 6 shows that TSA is more efficient than other GTTA baselines.
+A detailed discussion of computational efficiency is provided in Section 5.2.
+D.4 Hyperparameter sensitivity
+Hyperparameter analysis on $\rho_{1}$.
+TSA employs neighborhood alignment based on the uncertainty of target pseudo labels.
+To evaluate the sensitivity of the hyperparameter $\rho_{1}$ (as defined in Section 4.1 by the entropy threshold $-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{1}\cdot\ln(|{\mathcal{Y}}|)$), we vary $\rho_{1}$ within the range of $[-0.1,0.1]$ centered around the value selected on the validation set.
+We perform evaluations on both synthetic and real-world datasets, using the best-performing variants: TSA-TENT, TSA-T3A and TSA-LAME for each case.
+TSA-TENT and TSA-T3A are evaluated with the GraphSAGE backbone, while TSA-LAME is evaluated with the GPRGNN backbone.
+Overall, Figure 6 shows that TSA is more robust to variations in $\rho_{1}$ on synthetic datasets, but more sensitive on real-world datasets due to the greater task complexity.
+Nevertheless, all tested values of $\rho_{1}$ outperform the corresponding non-graph TTA baselines.
+Hyperparameter analysis on $\rho_{2}$.
+When optimizing $\alpha$ using the cross-entropy loss in Eq.
+5, real-world datasets often produce unreliable hard pseudo-labels.
+To address this, we adopt the same strategy used for selecting $\rho_{1}$.
+Only hard pseudo-labels whose entropy satisfies $-\sum_{i\in{\mathcal{Y}}}[\hat{y}]_{i}\ln([\hat{y}]_{i})\leq\rho_{2}\cdot\ln(|{\mathcal{Y}}|)$ are used to optimize $\alpha$.
+We select $\rho_{2}$ from $\{0.1,1.0\}$ for real-world datasets and fix $\rho_{2}=1$ for synthetic datasets.
+To further analyze the effect of $\rho_{2}$, we vary it within the range $[-0.1,0.1]$ centered around the value selected on the validation set.
+Figure 7 shows that  TSA is robust to variations in $\rho_{2}$, with all perturbations outperforming the corresponding non-graph TTA baselines.
+(a) CSBM (Struct. Shift) (b) MAG (US $\rightarrow$ FR) (c) Pileup (PU10$\rightarrow$30)
+Refer to caption Refer to caption Refer to caption
+Figure 6: Hyperparameter sensitivity analysis of $\rho_{1}$.
+(a) MAG (US $\rightarrow$ FR) (b) Pileup (PU10$\rightarrow$30)
+Refer to caption Refer to caption
+Figure 7: Hyperparameter sensitivity analysis of $\rho_{2}$ on real-world datasets.
+Table 7: GPRGNN results on MAG (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
 […]
@@
 GTrans 26.04$\pm$1.41 26.96$\pm$0.72 38.71$\pm$0.96 23.07$\pm$1.08 20.80$\pm$0.70 27.49$\pm$0.46 10.97$\pm$0.56 17.41$\pm$0.78 12.58$\pm$0.75 9.08$\pm$0.41
+HomoTTT 26.08$\pm$1.35 28.12$\pm$0.85 39.13$\pm$1.10 24.16$\pm$1.17 21.85$\pm$0.64 25.50$\pm$0.39 9.64$\pm$0.41 15.95$\pm$0.62 11.10$\pm$0.61 8.08$\pm$0.31
 SOGA 34.70$\pm$0.71 35.18$\pm$0.33 48.47$\pm$0.30 33.63$\pm$0.73 27.98$\pm$0.42 35.16$\pm$0.57 18.50$\pm$0.45 28.35$\pm$1.04 20.11$\pm$0.71 14.73$\pm$0.58
+ActMAD 25.14$\pm$1.41 26.77$\pm$0.69 37.97$\pm$1.09 22.30$\pm$0.99 20.52$\pm$0.61 26.89$\pm$0.57 10.35$\pm$0.52 16.90$\pm$0.85 12.03$\pm$0.77 8.71$\pm$0.42
 TENT 13.79$\pm$0.72 25.67$\pm$0.50 29.13$\pm$0.80 18.00$\pm$0.14 20.85$\pm$0.26 16.29$\pm$0.82 7.59$\pm$0.73 13.81$\pm$0.74 7.82$\pm$0.74 5.94$\pm$0.34
-LAME 24.54$\pm$2.55 22.50$\pm$0.85 35.97$\pm$1.42 17.61$\pm$1.46 15.71$\pm$0.92 28.02$\pm$0.55 9.08$\pm$0.61 15.58$\pm$1.02 10.87$\pm$0.70 8.01$\pm$0.38
+LAME 24.54$\pm$2.55 22.50$\pm$0.85 35.93$\pm$1.42 17.61$\pm$1.46 15.71$\pm$0.92 28.02$\pm$0.55 9.08$\pm$0.61 15.58$\pm$1.02 10.87$\pm$0.70 8.01$\pm$0.38
 T3A 40.58$\pm$0.23 39.00$\pm$0.85 48.96$\pm$1.00 46.13$\pm$1.56 32.35$\pm$1.29 38.92$\pm$2.23 38.07$\pm$1.45 46.23$\pm$1.73 42.82$\pm$1.52 28.65$\pm$0.65
-AdaRC-TENT 13.90$\pm$0.71 25.87$\pm$0.47 29.30$\pm$0.82 18.14$\pm$0.19 21.03$\pm$0.26 16.38$\pm$0.80 7.62$\pm$0.73 13.84$\pm$0.74 7.85$\pm$0.81 5.97$\pm$0.36
-AdaRC-LAME 8.93$\pm$4.76 8.18$\pm$0.78 15.46$\pm$1.34 6.25$\pm$1.43 6.69$\pm$0.81 1.35$\pm$2.69 17.60$\pm$8.89 33.42$\pm$2.41 18.81$\pm$1.42 4.23$\pm$8.02
-AdaRC-T3A 22.34$\pm$18.21 32.18$\pm$1.70 42.85$\pm$1.15 34.77$\pm$2.62 25.75$\pm$0.62 45.27$\pm$2.57 33.01$\pm$1.17 42.08$\pm$0.82 38.57$\pm$1.50 29.30$\pm$0.47
-TSA-TENT 14.25$\pm$0.68 25.76$\pm$0.51 29.30$\pm$0.79 18.30$\pm$0.26 21.20$\pm$0.20 16.57$\pm$0.78 7.69$\pm$0.72 14.03$\pm$0.70 8.07$\pm$0.79 6.03$\pm$0.39
-TSA-LAME 26.33$\pm$2.25 22.76$\pm$1.44 35.57$\pm$1.43 18.15$\pm$2.08 16.94$\pm$1.06 29.70$\pm$0.45 11.04$\pm$0.61 17.90$\pm$0.89 12.46$\pm$0.43 9.32$\pm$0.51
-TSA-T3A 40.81$\pm$0.23 39.13$\pm$0.77 49.16$\pm$0.85 47.05$\pm$1.74 33.27$\pm$1.21 41.06$\pm$3.34 38.34$\pm$1.25 46.38$\pm$1.83 43.64$\pm$0.85 28.71$\pm$0.57
-B.2 Results of Ablation Study
-Table 6: Ablation Study on TSA.
+Matcha-TENT 13.90$\pm$0.71 25.87$\pm$0.47 29.30$\pm$0.82 18.14$\pm$0.19 21.03$\pm$0.26 16.38$\pm$0.80 7.62$\pm$0.73 13.84$\pm$0.74 7.85$\pm$0.81 5.97$\pm$0.36
+Matcha-LAME 8.93$\pm$4.76 8.18$\pm$0.78 15.46$\pm$1.34 6.25$\pm$1.43 6.69$\pm$0.81 1.35$\pm$2.69 17.60$\pm$8.89 33.42$\pm$2.41 18.81$\pm$1.42 4.23$\pm$8.02
+Matcha-T3A 22.34$\pm$18.21 32.18$\pm$1.70 42.85$\pm$1.15 34.77$\pm$2.62 25.75$\pm$0.62 45.27$\pm$2.57 33.01$\pm$1.17 42.08$\pm$0.82 38.57$\pm$1.50 29.30$\pm$0.47
+TSA-TENT 14.51$\pm$0.66 25.86$\pm$0.54 29.47$\pm$0.73 18.76$\pm$0.39 21.48$\pm$0.23 16.96$\pm$0.74 7.89$\pm$0.72 14.32$\pm$0.70 8.35$\pm$0.81 6.16$\pm$0.38
+TSA-LAME 26.58$\pm$2.09 22.75$\pm$0.85 35.95$\pm$0.98 19.30$\pm$1.99 17.49$\pm$0.79 32.30$\pm$0.67 14.86$\pm$1.12 21.88$\pm$0.95 15.02$\pm$0.86 11.13$\pm$0.51
+TSA-T3A 40.77$\pm$0.15 39.03$\pm$0.86 49.38$\pm$0.89 47.81$\pm$1.81 32.43$\pm$1.24 40.94$\pm$2.82 38.12$\pm$1.42 46.26$\pm$1.81 43.70$\pm$1.33 29.07$\pm$0.46
+Table 8: GPRGNN results on Pileup (f1-scores). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
 […]
-TENT+TSA Base 27.30$\pm$1.61 32.84$\pm$0.78 40.82$\pm$0.99 32.53$\pm$0.91 28.62$\pm$0.63 27.89$\pm$0.98 16.22$\pm$1.17 24.87$\pm$0.56 22.05$\pm$0.84 14.20$\pm$0.71
-w/o Nbr. Align 27.18$\pm$1.70 32.75$\pm$0.69 40.82$\pm$0.99 32.26$\pm$0.95 28.41$\pm$0.59 27.53$\pm$0.92 16.22$\pm$1.17 24.87$\pm$0.56 22.05$\pm$0.84 13.95$\pm$0.76
-w/o SNR 27.00$\pm$1.21 32.76$\pm$0.75 40.80$\pm$0.91 32.52$\pm$0.90 28.57$\pm$0.59 27.70$\pm$0.88 15.84$\pm$0.85 24.62$\pm$0.48 21.37$\pm$0.73 14.11$\pm$0.60
-LAME+TSA Base 37.95$\pm$0.97 36.29$\pm$1.65 46.86$\pm$1.13 32.86$\pm$2.23 27.22$\pm$1.48 44.83$\pm$0.88 28.51$\pm$0.44 39.80$\pm$0.99 24.54$\pm$0.87 22.39$\pm$0.30
-w/o Nbr. Align 37.90$\pm$0.97 36.08$\pm$1.73 46.75$\pm$0.88 32.70$\pm$2.29 26.49$\pm$1.76 44.83$\pm$0.88 28.41$\pm$0.36 39.72$\pm$0.99 24.24$\pm$0.75 22.23$\pm$0.32
-w/o SNR 35.35$\pm$0.84 33.45$\pm$1.60 44.51$\pm$1.22 29.89$\pm$1.94 24.41$\pm$1.43 40.17$\pm$1.11 22.44$\pm$0.99 32.57$\pm$1.46 17.33$\pm$0.47 17.07$\pm$0.61
-T3A+TSA Base 41.65$\pm$0.99 47.01$\pm$2.08 51.65$\pm$0.90 46.61$\pm$0.88 43.45$\pm$0.81 48.09$\pm$0.60 39.18$\pm$1.87 46.50$\pm$0.25 43.70$\pm$1.38 30.89$\pm$2.13
-w/o Nbr. Align 41.64$\pm$1.10 46.37$\pm$2.11 50.32$\pm$1.11 46.44$\pm$0.80 42.09$\pm$1.26 46.47$\pm$1.27 39.12$\pm$1.85 46.23$\pm$0.61 43.60$\pm$1.21 30.92$\pm$2.03
-w/o SNR 41.59$\pm$0.94 45.54$\pm$1.81 51.55$\pm$0.95 46.51$\pm$0.94 41.35$\pm$1.09 48.11$\pm$0.62 38.71$\pm$1.00 46.29$\pm$0.26 43.12$\pm$0.76 29.90$\pm$1.35
-B.3 Visualization of Distribution Shift
-(a) Source (b) Feature Shift (c) Neighborhood Shift (d) Label Shift (e) SNR
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 4: The t-SNE visualization of a one-layer GSN under different distribution shifts: (top) the mean-pooled neighborhood hidden representations and (bottom) the node representations before passing through the classifier, positioned along its decision boundary.
-(a) indicates the source domain where the model is pretrained.
-(b), (c), and (d) stand for feature shift, neighborhood shift, and label shift.
-(e) represent the impact of SNR induced by degree shift.
-The color of the nodes represents the ground-truth labels.
-The legends in the bottom plots show the accuracy and recall scores for each class.
-The red contours in the bottom (d) highlight the decision boundary of the minority class.
-Appendix C Dataset Details
-C.1 Dataset Statistics
-Here we summarize the statistics of the four real-world graph datasets. The edges are counted twice in the edge list as they are undirected graphs.
+ERM 50.60$\pm$3.52 68.09$\pm$0.74 32.60$\pm$5.55 64.59$\pm$1.16 0.68$\pm$0.40 24.95$\pm$5.32 71.00$\pm$0.61 69.56$\pm$1.06
+GTrans 51.08$\pm$3.06 66.26$\pm$1.77 33.94$\pm$5.39 62.53$\pm$1.11 0.75$\pm$0.52 25.46$\pm$5.19 70.11$\pm$1.13 68.40$\pm$1.93
+HomoTTT 52.05$\pm$3.09 67.46$\pm$0.94 35.84$\pm$4.98 63.59$\pm$1.11 0.75$\pm$0.43 23.03$\pm$5.46 71.06$\pm$0.50 69.72$\pm$1.10
+SOGA 54.18$\pm$3.08 66.13$\pm$0.70 42.30$\pm$4.74 59.97$\pm$1.70 1.16$\pm$0.84 18.68$\pm$5.49 69.72$\pm$1.27 68.39$\pm$1.75
+ActMAD 58.80$\pm$1.58 62.22$\pm$0.95 49.22$\pm$2.21 56.04$\pm$2.24 2.16$\pm$0.75 17.81$\pm$6.45 70.40$\pm$0.94 69.95$\pm$1.02
+TENT 36.64$\pm$0.15 2.04$\pm$1.05 27.01$\pm$0.06 0.62$\pm$0.81 12.74$\pm$0.50 3.77$\pm$3.30 60.19$\pm$3.55 71.19$\pm$0.58
+LAME 46.33$\pm$4.67 67.61$\pm$0.96 22.58$\pm$9.42 64.58$\pm$1.00 0.00$\pm$0.00 20.94$\pm$7.27 70.72$\pm$0.74 69.67$\pm$0.88
+T3A 55.66$\pm$2.89 71.36$\pm$0.79 47.89$\pm$4.19 71.01$\pm$0.38 19.69$\pm$3.72 54.38$\pm$2.56 70.34$\pm$1.33 68.63$\pm$1.51
+Matcha-TENT 36.58$\pm$0.17 0.00$\pm$0.00 27.01$\pm$0.06 0.00$\pm$0.00 9.17$\pm$0.94 0.00$\pm$0.00 0.72$\pm$0.76 20.30$\pm$4.03
+Matcha-LAME 7.45$\pm$3.55 0.00$\pm$0.00 8.28$\pm$4.16 0.00$\pm$0.00 0.00$\pm$0.00 0.00$\pm$0.00 0.52$\pm$1.05 6.56$\pm$4.08
+Matcha-T3A 15.48$\pm$2.91 0.00$\pm$0.00 14.15$\pm$1.86 0.00$\pm$0.00 0.00$\pm$0.00 9.60$\pm$19.19 17.20$\pm$14.72 22.77$\pm$5.23
+TSA-TENT 39.90$\pm$1.18 5.07$\pm$10.13 27.78$\pm$0.48 0.59$\pm$0.75 20.31$\pm$0.65 3.65$\pm$3.22 67.97$\pm$1.60 71.60$\pm$0.18
+TSA-LAME 60.33$\pm$0.80 71.54$\pm$0.33 52.23$\pm$0.61 70.62$\pm$0.35 38.32$\pm$0.27 20.49$\pm$7.76 72.28$\pm$0.47 71.23$\pm$0.27
+TSA-T3A 59.40$\pm$1.70 72.94$\pm$0.18 53.13$\pm$0.34 72.63$\pm$0.23 36.39$\pm$1.48 62.31$\pm$0.28 71.30$\pm$1.23 69.70$\pm$1.04
+Table 9: GPRGNN results on Arxiv and DBLP/ACM (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+1950-2007 1950-2009 1950-2011 DBLP & ACM
+[…]
+ERM 45.62$\pm$2.30 40.35$\pm$4.46 49.21$\pm$1.36 44.78$\pm$2.70 56.49$\pm$0.97 54.14$\pm$1.31 67.34$\pm$1.21 56.47$\pm$2.84
+GTrans 45.25$\pm$2.25 40.51$\pm$4.31 48.56$\pm$1.23 44.57$\pm$2.55 55.81$\pm$0.99 53.59$\pm$1.34 59.61$\pm$0.98 59.27$\pm$1.79
+HomoTTT 45.49$\pm$2.64 40.03$\pm$4.80 48.10$\pm$1.17 42.33$\pm$2.34 56.27$\pm$1.03 53.25$\pm$1.24 67.11$\pm$1.03 57.18$\pm$2.54
+SOGA 48.40$\pm$2.29 46.01$\pm$4.00 52.59$\pm$1.38 50.64$\pm$1.63 58.61$\pm$1.15 57.68$\pm$0.88 67.00$\pm$1.03 55.42$\pm$3.03
+ActMAD 45.47$\pm$2.28 40.12$\pm$4.42 49.42$\pm$1.31 44.96$\pm$2.59 56.63$\pm$1.00 54.25$\pm$1.25 67.07$\pm$1.11 55.83$\pm$2.76
+TENT 44.73$\pm$2.63 40.58$\pm$3.44 48.47$\pm$0.94 42.37$\pm$1.52 57.15$\pm$0.98 52.45$\pm$1.75 61.56$\pm$0.77 56.84$\pm$1.04
+LAME 45.75$\pm$3.96 41.35$\pm$7.20 49.52$\pm$2.24 50.18$\pm$3.36 55.65$\pm$1.09 54.93$\pm$1.72 66.92$\pm$0.99 56.58$\pm$2.83
+T3A 32.73$\pm$1.63 31.97$\pm$1.75 36.99$\pm$2.01 34.10$\pm$1.32 40.53$\pm$2.88 39.04$\pm$3.86 67.21$\pm$1.31 65.32$\pm$1.69
+Matcha-TENT 43.89$\pm$2.77 39.30$\pm$3.63 48.01$\pm$0.90 41.91$\pm$1.66 56.28$\pm$1.30 51.27$\pm$1.64 64.43$\pm$0.90 78.61$\pm$1.38
+Matcha-LAME 32.93$\pm$3.30 24.13$\pm$5.06 37.41$\pm$2.06 24.16$\pm$2.92 48.29$\pm$3.27 44.74$\pm$4.20 67.83$\pm$1.71 70.39$\pm$2.15
+Matcha-T3A 24.81$\pm$2.32 19.06$\pm$3.09 25.85$\pm$3.54 19.60$\pm$4.56 26.90$\pm$1.55 18.74$\pm$4.37 67.59$\pm$1.87 72.91$\pm$2.69
+TSA-TENT 45.17$\pm$2.84 41.07$\pm$3.67 49.13$\pm$0.84 43.19$\pm$1.24 57.66$\pm$1.00 53.00$\pm$1.65 62.77$\pm$0.84 61.71$\pm$0.99
+TSA-LAME 49.75$\pm$4.12 48.67$\pm$7.74 53.82$\pm$0.53 55.18$\pm$1.21 57.67$\pm$0.98 57.14$\pm$1.83 67.71$\pm$1.10 65.20$\pm$1.70
+TSA-T3A 39.11$\pm$1.73 39.40$\pm$3.69 41.15$\pm$6.33 44.65$\pm$2.31 44.14$\pm$2.63 40.82$\pm$5.28 68.02$\pm$1.40 74.64$\pm$1.38
+Table 10: GCN results on MAG (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+[…]
+ERM 27.76$\pm$1.49 30.70$\pm$0.96 40.12$\pm$0.73 27.70$\pm$0.75 23.03$\pm$0.80 36.46$\pm$2.00 21.03$\pm$1.56 28.95$\pm$2.43 20.53$\pm$2.07 16.68$\pm$1.54
+GTrans 27.44$\pm$1.48 30.59$\pm$0.99 39.65$\pm$0.97 27.79$\pm$0.72 22.95$\pm$0.88 35.88$\pm$2.09 20.80$\pm$1.59 28.13$\pm$2.49 20.53$\pm$2.06 16.52$\pm$1.51
+HomoTTT 23.64$\pm$2.09 28.49$\pm$1.29 35.34$\pm$1.87 26.21$\pm$1.62 23.78$\pm$0.56 29.27$\pm$2.91 18.60$\pm$1.59 23.10$\pm$2.01 21.71$\pm$1.70 16.03$\pm$1.33
+SOGA 28.36$\pm$2.23 35.61$\pm$1.50 45.29$\pm$2.08 36.79$\pm$1.18 29.30$\pm$0.90 44.09$\pm$1.92 26.78$\pm$1.26 37.93$\pm$2.05 33.24$\pm$1.49 24.23$\pm$1.01
+ActMAD 28.11$\pm$1.34 31.25$\pm$0.90 40.71$\pm$0.65 28.61$\pm$0.63 23.82$\pm$0.74 36.07$\pm$2.01 20.73$\pm$1.50 28.76$\pm$2.33 20.92$\pm$2.04 16.65$\pm$1.46
+TENT 26.66$\pm$0.57 33.38$\pm$0.32 41.84$\pm$0.59 33.01$\pm$0.54 29.17$\pm$0.20 28.82$\pm$0.85 16.35$\pm$0.76 25.20$\pm$0.85 23.45$\pm$0.87 14.96$\pm$0.56
+LAME 28.20$\pm$3.24 30.90$\pm$1.77 42.25$\pm$0.97 27.20$\pm$2.17 20.83$\pm$1.08 35.89$\pm$2.53 19.31$\pm$2.11 30.01$\pm$4.10 17.46$\pm$2.39 14.24$\pm$2.02
+T3A 41.53$\pm$0.93 45.41$\pm$1.76 51.11$\pm$1.18 48.02$\pm$1.99 40.09$\pm$1.53 43.72$\pm$2.31 39.68$\pm$2.25 47.31$\pm$0.84 45.93$\pm$1.14 30.95$\pm$1.20
+TSA-TENT 27.46$\pm$0.53 33.89$\pm$0.31 42.34$\pm$0.60 33.72$\pm$0.53 29.59$\pm$0.21 29.33$\pm$0.86 16.82$\pm$0.75 25.90$\pm$0.89 24.03$\pm$0.81 15.32$\pm$0.56
+TSA-LAME 28.64$\pm$3.24 31.01$\pm$1.91 42.35$\pm$0.86 27.22$\pm$1.87 22.27$\pm$0.52 39.64$\pm$2.08 22.96$\pm$1.62 33.94$\pm$3.14 20.08$\pm$1.93 17.36$\pm$1.91
+TSA-T3A 41.61$\pm$0.95 46.13$\pm$1.83 51.81$\pm$1.22 48.87$\pm$1.43 42.98$\pm$1.90 45.48$\pm$2.69 40.11$\pm$2.74 47.45$\pm$0.42 45.95$\pm$1.17 32.03$\pm$1.30
+Table 11: GCN results on Pileup (f1-scores). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+[…]
+ERM 8.30$\pm$1.79 29.20$\pm$4.57 2.27$\pm$0.70 26.61$\pm$6.79 0.33$\pm$0.13 16.08$\pm$10.46 23.71$\pm$2.49 26.73$\pm$4.79
+GTrans 11.97$\pm$2.67 28.88$\pm$4.71 3.90$\pm$1.26 27.20$\pm$6.88 0.91$\pm$0.29 16.98$\pm$11.06 24.62$\pm$2.38 28.88$\pm$5.14
+HomoTTT 8.94$\pm$2.05 31.21$\pm$4.79 2.54$\pm$0.81 27.59$\pm$6.18 0.55$\pm$0.18 16.77$\pm$11.23 24.84$\pm$3.27 27.48$\pm$4.95
+SOGA 5.81$\pm$2.20 35.21$\pm$4.00 1.00$\pm$0.44 27.19$\pm$6.76 0.54$\pm$0.20 19.17$\pm$11.32 25.38$\pm$1.66 28.06$\pm$5.06
+ActMAD 7.20$\pm$2.22 30.45$\pm$4.13 1.15$\pm$0.41 24.35$\pm$6.69 0.31$\pm$0.14 16.67$\pm$9.32 23.77$\pm$1.78 27.36$\pm$4.75
+TENT 2.45$\pm$0.93 35.99$\pm$3.12 1.63$\pm$0.75 26.20$\pm$7.51 0.09$\pm$0.05 13.84$\pm$8.89 19.96$\pm$4.06 29.58$\pm$4.46
+LAME 7.48$\pm$2.17 30.91$\pm$5.58 1.46$\pm$0.75 24.97$\pm$9.02 0.06$\pm$0.06 15.30$\pm$10.74 21.50$\pm$5.24 25.72$\pm$6.51
+T3A 41.55$\pm$13.17 62.80$\pm$2.48 29.92$\pm$7.29 56.32$\pm$14.68 4.45$\pm$2.48 33.26$\pm$4.35 60.15$\pm$10.72 62.77$\pm$8.69
+TSA-TENT 2.96$\pm$1.12 38.45$\pm$3.83 2.17$\pm$0.82 26.85$\pm$7.89 0.24$\pm$0.16 15.20$\pm$10.17 24.22$\pm$3.88 33.47$\pm$4.43
+TSA-LAME 12.71$\pm$2.29 32.37$\pm$8.02 4.14$\pm$1.48 27.14$\pm$10.60 0.38$\pm$0.19 16.53$\pm$12.07 29.29$\pm$2.32 27.72$\pm$4.22
+TSA-T3A 43.86$\pm$13.14 64.00$\pm$1.56 33.31$\pm$8.50 60.22$\pm$8.38 7.12$\pm$4.19 35.72$\pm$4.56 62.03$\pm$8.01 64.02$\pm$6.87
+Table 12: GCN results on Arxiv and DBLP/ACM (accuracy). Bold indicates improvements in comparison to the corresponding non-graph TTA baselines. Underline indicates the best model.
+1950-2007 1950-2009 1950-2011 DBLP & ACM
+[…]
+ERM 49.25$\pm$1.58 50.64$\pm$3.11 52.61$\pm$2.67 51.92$\pm$4.12 58.83$\pm$0.83 57.73$\pm$1.27 62.56$\pm$3.32 47.59$\pm$1.92
+GTrans 49.10$\pm$1.68 50.57$\pm$3.10 52.50$\pm$2.71 51.87$\pm$4.26 58.72$\pm$0.89 57.75$\pm$1.31 62.03$\pm$2.31 51.65$\pm$3.09
+HomoTTT 46.22$\pm$1.54 46.11$\pm$2.79 48.96$\pm$3.13 46.83$\pm$5.00 55.10$\pm$1.63 53.18$\pm$2.52 63.21$\pm$3.31 48.45$\pm$1.68
+SOGA 49.82$\pm$2.38 52.36$\pm$2.54 52.92$\pm$1.08 54.45$\pm$1.15 57.93$\pm$1.71 58.66$\pm$2.02 64.22$\pm$2.40 53.29$\pm$2.08
+ActMAD 49.24$\pm$1.57 50.68$\pm$2.94 52.71$\pm$2.64 52.10$\pm$3.98 58.83$\pm$0.86 57.85$\pm$1.25 62.68$\pm$3.36 47.58$\pm$1.89
+TENT 48.59$\pm$1.41 49.90$\pm$1.15 53.39$\pm$1.82 53.45$\pm$2.11 58.72$\pm$0.91 58.23$\pm$0.95 60.80$\pm$2.51 48.11$\pm$1.43
+LAME 49.77$\pm$2.50 50.97$\pm$4.60 53.35$\pm$2.72 52.70$\pm$4.25 58.85$\pm$0.57 57.05$\pm$1.57 62.51$\pm$3.33 47.64$\pm$1.70
+T3A 43.37$\pm$1.38 41.32$\pm$3.13 49.95$\pm$3.76 47.11$\pm$6.73 55.56$\pm$1.16 53.57$\pm$1.94 62.60$\pm$3.65 49.65$\pm$1.83
+TSA-TENT 49.63$\pm$1.26 50.78$\pm$0.74 53.78$\pm$1.57 53.67$\pm$1.85 58.90$\pm$0.94 58.27$\pm$0.95 61.31$\pm$2.39 53.14$\pm$2.01
+TSA-LAME 50.95$\pm$2.05 53.04$\pm$2.86 54.49$\pm$1.81 54.13$\pm$3.35 59.26$\pm$0.70 57.55$\pm$1.30 63.92$\pm$2.76 55.19$\pm$4.50
+TSA-T3A 45.37$\pm$0.76 45.46$\pm$1.92 51.57$\pm$2.73 49.70$\pm$5.44 56.14$\pm$1.02 54.84$\pm$1.17 63.31$\pm$3.50 54.53$\pm$3.07
+D.5 Results of More GNN Architectures
+TSA is Model-Agonistic.
+TSA employs neighborhood alignment and SNR adjustment on edge weights and do not need any assumption on the GNN architecture such as hop-aggregation parameters.
+We demonstrate that  TSA is applicable to common GNN architectures such as GraphSAGE, GPRGNN, and GCN.
+While different backbones yield varying ERM results, we observe that  TSA consistently improves model performance.
+Comparison with Baselines.
+TSA exhibits strong outcomes on real-world datasets and with both GPRGNN and GCN backbones.
+In all cases, the best results are obtained by GTTA methods (underlined across Table712), highlighting the importance of addressing structure shift in graph data.
+For GPRGNN,  TSA consistently outperforms GTrans and HomoTTT and surpasses SOGA by an average of 10.5% and Matcha by an average of 19.5%.
+For GCN,  TSA similarly outperforms GTrans and HomoTTT, and exceeds SOGA by an average of 13.1%.
+Furthermore, TSA shows prevailing gains over the corresponding non-graph TTA baselines (bold across Table712).
+While Matcha sometimes demonstrates strong performance, we frequently observe unstable results that may even fall below ERM.
+This drop is attributed to the reliance of their proposed PIC loss on minimizing intra-variance, which depends on the quality of node hidden representations in the target domain.
+Under label shift, node representations can easily degrade, leading to instability in Matcha’s performance.
+Appendix E Dataset Details
+E.1 Dataset Statistics
+Here we summarize the statistics of the four real-world graph datasets in Table LABEL:table:arxiv_datastats, LABEL:table:mag_datastats, and LABEL:table:pileup_datastats.
+The edges are counted twice in the edge list as they are undirected graphs.
 Table 13: Arxiv and ACM/DBLP dataset statistics
@@
 Table 15: Pileup dataset statistics
-gg-10 qq-10 gg-30 qq-30 gg-50 gg-140
-Nodes $18611$ $17242$ $41390$ $38929$ $60054$ $154750$
-Edges $53725$ $42769$ $173392$ $150026$ $341930$ $2081229$
-Features $28$ $28$ $28$ $28$ $28$ $28$
-Classes $2$ $2$ $2$ $2$ $2$ $2$
-C.2 Dataset Shift Statistics
+gg-10 qq-10 gg-30 gg-50 gg-140
+Nodes $18611$ $17242$ $41390$ $60054$ $154750$
+Edges $53725$ $42769$ $173392$ $341930$ $2081229$
+Features $28$ $28$ $28$ $28$ $28$
+Classes $2$ $2$ $2$ $2$ $2$
+E.2 Dataset Shift Statistics
 Below we provide two metrics indicating the neighborhood and label shift in real-world dataset.
-We measure the neighborhood shift by computing the weighted average of the total variation distance between of the neighborhood distribution:
-[…]
+We measure the CSS by computing the weighted average of the total variation distance between of the neighborhood distribution:
+$\text{CSS}\coloneqq\frac{1}{2}\sum_{j\in{\mathcal{Y}}}\sum_{k\in{\mathcal{Y}}}\mathbb{P}^{\mathcal{T}}(Y_{u}=j)\lvert\mathbb{P}^{\mathcal{S}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})-\mathbb{P}^{\mathcal{T}}(Y_{v}=k|Y_{u}=j,v\in\mathcal{N}_{u})\rvert$
 The label shift is measured as the total variation distance between the label distribution, which corresponds to the first term in Theorem 3.3.
-[…]
+Label $\text{Label Shift}\coloneqq\frac{1}{2}\sum_{j\in{\mathcal{Y}}}\lvert\mathbb{P}^{\mathcal{S}}(Y_{u}=j)-\mathbb{P}^{\mathcal{T}}(Y_{u}=j)\rvert$
 Table 16: MAG dataset shift metrics
-$US\rightarrow CN$ $US\rightarrow DE$ $US\rightarrow JP$ $US\rightarrow RU$ $US\rightarrow FR$ $CN\rightarrow US$ $CN\rightarrow DE$ $CN\rightarrow JP$ $CN\rightarrow RU$ $CN\rightarrow FR$ Nbr.
-Shift $0.2137$ $0.1902$ $0.1476$ $0.2855$ $0.2238$ $0.1946$ $0.2053$ $0.1361$ $0.2025$ $0.2039$ Label Shift $0.2734$ $0.1498$ $0.1699$ $0.3856$ $0.1706$ $0.2734$ $0.2691$ $0.1522$ $0.2453$ $0.2256$
+$US\rightarrow CN$ $US\rightarrow DE$ $US\rightarrow JP$ $US\rightarrow RU$ $US\rightarrow FR$ $CN\rightarrow US$ $CN\rightarrow DE$ $CN\rightarrow JP$ $CN\rightarrow RU$ $CN\rightarrow FR$ CSS $0.2137$ $0.1902$ $0.1476$ $0.2855$ $0.2238$ $0.1946$ $0.2053$ $0.1361$ $0.2025$ $0.2039$ Label Shift $0.2734$ $0.1498$ $0.1699$ $0.3856$ $0.1706$ $0.2734$ $0.2691$ $0.1522$ $0.2453$ $0.2256$
 Table 17: Pileup dataset shift metrics
-Pileup Conditions Physical Processes Domains $\text{PU}10\rightarrow 30$ $\text{PU}30\rightarrow 10$ PU1050105010$\rightarrow$ 5010 → 50 PU5010501050$\rightarrow$ 1050 → 10 PU301403014030$\rightarrow$ 14030 → 140 PU1403014030140$\rightarrow$ 30140 → 30 $gg\rightarrow qq$ $qq\rightarrow gg$ Nbr.
-Shift $0.1493$ $0.1709$ $0.2039$ $0.2595$ $0.1293$ $0.1635$ $0.0214$ $0.0205$ Label Shift $0.2425$ $0.2425$ $0.2982$ $0.2982$ $0.1377$ $0.1377$ $0.0341$ $0.0348$
+Pileup Conditions Physical Processes Domains $\text{PU}10\rightarrow 30$ $\text{PU}30\rightarrow 10$ PU1050105010$\rightarrow$ 5010 → 50 PU5010501050$\rightarrow$ 1050 → 10 PU301403014030$\rightarrow$ 14030 → 140 PU1403014030140$\rightarrow$ 30140 → 30 $gg\rightarrow qq$ $qq\rightarrow gg$ CSS $0.1493$ $0.1709$ $0.2039$ $0.2595$ $0.1293$ $0.1635$ $0.0214$ $0.0205$ Label Shift $0.2425$ $0.2425$ $0.2982$ $0.2982$ $0.1377$ $0.1377$ $0.0341$ $0.0348$
 Table 18: Real dataset shift metrics
-1950-2007 1950-2009 1950-2011 DBLP/ACM Domains $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $D\rightarrow A$ $A\rightarrow D$ Nbr.
-Shift $0.3005$ $0.3450$ $0.2583$ $0.3120$ $0.1833$ $0.2567$ $0.1573$ $0.2227$ Label Shift $0.2938$ $0.4396$ $0.2990$ $0.4552$ $0.2853$ $0.4438$ $0.3434$ $0.3435$
-Appendix D Detailed Experimental Setup
-D.1 Datasets Setup
-For Arxiv, MAG, and DBLP/ACM we follow the setting of (Liu et al., 2024). Hence we discuss the setting of synthetic dataset and pileup in the following:
+1950-2007 1950-2009 1950-2011 DBLP/ACM Domains $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $2014-2016$ $2016-2018$ $D\rightarrow A$ $A\rightarrow D$ CSS $0.3005$ $0.3450$ $0.2583$ $0.3120$ $0.1833$ $0.2567$ $0.1573$ $0.2227$ Label Shift $0.2938$ $0.4396$ $0.2990$ $0.4552$ $0.2853$ $0.4438$ $0.3434$ $0.3435$
+Appendix F Detailed Experimental Setup
+F.1 Datasets Setup
+For Arxiv, MAG, and DBLP/ACM we follow the setting of [59]. Hence we discuss the setting of synthetic dataset and pileup in the following:
 Synthetic Dataset.
-The synthetic dataset is generated by the contextual stochastic block model (CSBM) (Deshpande et al., 2018).
+The synthetic dataset is generated by the contextual stochastic block model (CSBM) [60].
 The generated graph contains 6000 nodes and 3 classes.
@@
 Under the study of different PU levels we use the data from signal $gg$.
-Compared to the Pileup datasets used in the (Liu et al., 2024), we explicitly label all types of particles to better align the neighborhood distribution.
+Compared to the Pileup datasets used in the [59], we explicitly label all types of particles to better align the neighborhood distribution.
 Note that the charged particle can achieve all most perfect recall as their label information is encoded in their feature, hence the four-class classification still reduces into a binary classification between LC neutral particles and OC neutral particles.
 For the study of different data generation process, we use PU10 from signal $gg\rightarrow qq$ and $qq\rightarrow gg$.
-D.2 Pretraining Setup
-Model architecture and Pretraining
+F.2 Pretraining Setup
+Model Architecture and Pretraining.
 We use GraphSAGE as the model backbone, with a 3-layer mean pooling GNN as the encoder and a 2-layer MLP with a batch normalization layer as the classifier.
@@
 For the GPRGNN backbone, we follow the same configuration but increase the number of encoder layers to 5.
+For the GCN backbone, we use a 3-layer GCN as the feature encoder, followed by the same 2-layer MLP classifier as in the other backbones.
 All models are trained for 400 epochs with the learning rate set to 0.003.
-A learning rate scheduler is applied with a decay rate of 0.9.
+We use Adam [75] as the optimizer for all models, and apply a learning rate scheduler with a decay rate of 0.9.
 All experiments are repeated 5 times under different initializations and data splits to ensure consistency.
-D.3 Evaluation and Metric
+Hardware and Frameworks.
+All experiments are conducted on a server with a single NVIDIA RTX 6000 GPU and an AMD EPYC 7763 64-Core Processor.
+We use PyTorch 2.2.2 [76] with CUDA 12.1 as the framework.
+For graph neural networks, we use PyTorch Geometric 2.5.2 [77].
+F.3 Evaluation and Metric
 The source graph is splitted into train/val/test 60/20/20 during the pretraining stage.
@@
 We use the the labeled nodes in the target graph to do hyperparameter tuning and select the model with the optimal validation score.
-We follow the metrics in (Liu et al., 2024) for evaluation.
+We follow the metrics in [59] for evaluation.
 MAG, Arxiv, DBLP/ACM, and synthetic datasets are evaluated with accuracy.
 For the MAG datasets, only the top 19 classes are evaluated.
-Pileup is evaluated with f1 score.
-D.4 Hyperparameter Tuning
+For the Pileup datasets, we evaluate the performance with f1 score.
+F.4 Hyperparameter Tuning
 Below we introduce the search space of the hyperparameters. LAME is a parameter-free approach so we do not tune it.
@@
 We observe that given the same domain shifts, the optimal hyperparameters of $\rho_{1}$ may differ due to different performance of the boundary refinement approaches (TENT, T3A, and LAME).
-  • AdaRC introduces two hyperparameters learning rate $lr$ and the number of epochs $T$. Based on their hyperparameter study, We select $lr$ from $\{1,5,10\}$ and $T$ from $\{1,10,100\}$.
+  • Matcha introduces two hyperparameters learning rate $lr$ and the number of epochs $T$. Based on their hyperparameter study, We select $lr$ from $\{0.1,1,5,10\}$ and $T$ from $\{2,10\}$.
   • GTrans introduces four hyperparameters learning rate of feature adaptation $lr_{f}$, learning rate of structure adaptation $lr_{A}$, the number of epochs $T$, and the budget to update.
 We select $lr_{f}$ from $\{0.001,0.01\}$, $lr_{A}=\{0.01,0.1,0.5\}$ from $\{0.01,0.05\}$, $T$ from $\{5,10\}$, and budget from $\{0.01,0.05,0.1\}$.
-  • SOGA introduces one hyperparameters $lr$ for one epoch update. We select $lr$ from $\{0.001,0.01\}$.
-  • TENT introduces one hyperparameters $lr$ for one epoch update. We select $lr$ from $\{0.001,0.01,0.05\}$.
-  • T3A introduces one hyperparameters $M$ for deciding the number of supports to restore. We select $M$ from $\{5,20,50,100\}$
+  • HomoTTT introduces two hyperparameter $lr$ and $k0$.
+The hyperparameter $k0$ is a scaling factor for the model selection score.
+We present the results updated by one epoch.
+We select $lr$ from $\{0.001,0.01\}$ and $k0$ from $\{0.001,0.01\}$.
+  • SOGA introduces one hyperparameter $lr$. We present the results updated by one epoch and select $lr$ from $\{0.001,0.01\}$.
+  • TENT introduces one hyperparameter $lr$. We present the results updated by one epoch and select $lr$ from $\{0.001,0.01,0.05\}$.
+  • T3A introduces one hyperparameter $M$ for deciding the number of supports to restore. We select $M$ from $\{5,20,50,100\}$
+  • ActMAD introduces one hyperparameter $lr$.
+We present the results updated by one epoch and select $lr$ from $\{0.001,0.01\}$.
+We adapt ActMAD to GNNs by aligning the outputs of each GNN layer before the classifier.